### v1

# Understanding Warmup-Stable-Decay Learning Rates: A River Valley Loss Landscape Perspective

Abstract
Training language models currently requires pre-determining a fixed compute budget because the typical cosine learning rate schedule depends on the total number of steps. In contrast, the Warmup-Stable-Decay (WSD) schedule uses a constant learning rate to produce a main branch of iterates that can in principle continue indefinitely without a pre-specified compute budget. Then, given any compute budget, one can branch out from the main branch at a proper at any time with a rapidly decaying learning rate to produce a strong model. Empirically, WSD generates an intriguing, non-traditional loss curve: the loss remains elevated during the stable phase but sharply declines during the decay phase. Towards explaining this phenomenon, we conjecture that pretraining loss exhibits a river valley landscape, which resembles a deep valley with a river at its bottom. Under this assumption, we show that during the stable phase, the iterate undergoes large oscillations due to the high learning rate, yet it progresses swiftly along the river. During the decay phase, the rapidly dropping learning rate minimizes the iterate’s oscillations, moving it closer to the river and revealing true optimization progress. Therefore, the sustained high learning rate phase and fast decaying phase are responsible for progress in the river and the mountain directions, respectively, and are both critical. Our analysis predicts phenomenons consistent with empirical observations and shows that this landscape can naturally emerge from pretraining on a simple bi-gram dataset. Inspired by the theory, we introduce WSD-S, a variant of WSD that reuses previous checkpoints’ decay phases and keeps only one main branch, where we resume from a decayed checkpoint. WSD-S empirically outperforms WSD and Cyclic-Cosine in obtaining multiple pretrained language model checkpoints across various compute budgets in a single run for parameters scaling from 0.1B to 1.2B.
Kaiyue Wen
Stanford University
kaiyuew@stanford.edu
Zhiyuan Li
TTIC
zhiyuanli@ttic.edu
Jason Wang
Stanford University
jsywang@stanford.edu
David Hall
Stanford University
dlwh@cs.stanford.edu
Percy Liang
Stanford University
pliang@cs.stanford.edu
Tengyu Ma
Stanford University
tengyuma@cs.stanford.edu
1 Introduction
Pre-training large language models (LLMs) typically involves following a learning rate schedule that decreases over a pre-determined number of steps, such as a cosine schedule (Loshchilov & Hutter, 2017; Touvron et al., 2023), where the learning rate starts high and gradually decreases in a smooth curve following the shape of a cosine function. This inflexible approach makes it difficult to adapt to additional compute or data, as the learning rate schedule for all the data is not a natural continuation of the schedule used with past data. Additionally, fitting scaling laws is costly because each compute budget requires retraining to adjust the learning rate schedule (Hoffmann et al., 2022).
In contrast to the cosine learning rate, recent work Hu et al. (2024) introduces the warmup-stable-decay (WSD) schedule, which does not require committing to a pre-specified total compute budget. After a standard warm-up period, the WSD schedule maintains a main “branch” using a constant learning rate indefinitely and branch off using a fast-decaying learning rate schedule to obtain intermediate checkpoints (see the second row of Figure 2(b)). Using the WSD schedule, one can continue training from a checkpoint in the main branch by resuming with the same constant learning rate and can obtain training losses for multiple compute budgets with a single run.
Empirically, the WSD schedule produces a non-traditional loss curve (see Figure 1): during the constant learning rate phase, the loss remains higher than the loss using other schedules like the cosine schedule; but during the decay phase, it drops sharply, often leading to a better final performance compared to the cosine schedule. This raises the main question the paper aims to address:
Figure 1: The Distinctive Loss Curve produced by WSD. A constant learning rate phase, characterized by slow loss improvements, eventually leads to better validation loss after learning rate decay.
Why does WSD work, especially with such a non-traditional loss curve? Specifically, why does a constant learning rate phase, characterized by slow loss improvements, eventually lead to superior performance?
(a) River Valley Landscape
(b) Learning rate schedules
Figure 2: We theoretically analyze the Warmup-Stable-Decay (WSD) schedule and demonstrate a river valley loss lanscape model to explain its effectiveness (demonstrated in Figure 2(a)). The stable phase adopts a large learning rate and the iterate will progress along the river while oscillating between the sharp hillsides. Due to the large oscillation caused by the large learning rate, the run will potentially show a higher loss compared to a run using smaller learning rate in this phase. During the decay phase, the learning rate is dropped rapidly to ease the oscillation of the iterates, driving it closer to the river, revealing the optimization progress. Based on our theory, we propose WSD-Simplified (WSD-S), an effective simplification of the WSD schedule in continual learning, where we start directly using a high learning rate from previous intermediate checkpoints. We visualize the learning rate schedule in Figure 2(b). The arrow in the second row of Figure 2(b) indicates WSD reinitializes the checkpoint from the last checkpoint from the constant learning rate phase instead.
The first contribution of this paper is a theoretical framework to explain the underlying mechanism of WSD. We characterize a type of loss landscape, called the river valley landscape (Definition 3.1), and theoretically show that WSD has superior performance on such loss landscapes. We show that the river valley landscape can provide multiple theoretical prediction matching the empirical observations and hence can serve as a useful conceptual picture for understanding the pretraining optimization process.
As the name suggests, a river valley landscape intuitively features steeply sloping hillsides with a river winding through the bottom of the gorge (see Figure 2(a)). During the stochastic gradient-based optimization process, the iterate bounces between the hillsides as it slowly and implicitly progresses along the river direction. The loss in this landscape can be decomposed into two components: the river component, which represents the primary loss along the river at the bottom of the hills, and the hill component, which accounts for the additional loss caused by deviations in height from the river’s course. Progress is determined primarily by the river component in the long run. We demonstrate that when the loss function exhibits this type of landscape, a learning rate schedule should satisfy the following two key properties to effectively minimize the loss.
  1. 1.
    Sustained high learning rate. It is advantageous to maintain a large learning rate for as long as possible during training, even at the cost of less reduction in the loss. A large learning rate yields larger bouncing due to the stochasticity of the gradient, increasing the hill component of the loss, but it also makes faster progress in the river direction. In contrast, a small learning rate results in less bouncing, keeping the iterate close to the river, but progress along the river direction is slower. Therefore, a larger learning rate leads to faster fundamental progress minimizing the river component, which is obscured by the oscillation in the hill component. This progress will be revealed by the decay phase discussed below.
  2. 2.
    Final low learning rate. As training nears completion, it becomes essential to reduce the learning rate. This decay minimizes the oscillations in the mountain direction to decrease the hill component and ensures that the iterates converge to a point close to the river, which has a lower loss than any nearby points up the mountain.
In Section 3, we provide formal theoretical statements analyzing the trajectories of (stochastic) gradient descent on the river valley landscape, fleshing out the intuitions above. Among our synthetic and real-world studies supporting the river valley landscape hypothesis, an intriguing observation in language model pretraining is that the loss on the linear interpolation of two checkpoints in the stable phase exhibits a convex and unimodal shape, resembling a valley, whereas between two checkpoints in the decay phase, the loss shows a smooth monotone decay.
All the theoretical results above assume a river valley landscape. How likely does the next-token prediction loss follow this pattern, and why? We hypothesize the river valley landscape can naturally arise from the heterogeneity in the stochasticity of different tokens: highly deterministic tokens (which often involve facts and knowledge) contribute to the "river" direction, while uncertain tokens (which often involve flexibility and ambiguity in the language) create the steep hillsides. We demonstrate this insight by showing in Section 4 that under a bigram toy model, indeed the loss has a river valley landscape, and empirically, most properties of the loss curves under various learning rates on the real datasets are still seen in this toy model. We further show that the stable learning rate phase learns the deterministic tokens whereas the decay phase learns better the stochastic tokens.
Finally, we discuss the limitations of WSD within our framework and how to go beyond it. In WSD, once an intermediate checkpoint is reached, the model and optimizer are rolled back to the end of the stable phase before resuming training with a constant learning rate. However, our theory suggests that the decay phase also contributes to progress along the river direction, so discarding this part of the progress is unnecessary. Moreover, when an organization pretrains an LLM with WSD, it needs to release the last checkpoint in a stable-phase to facilitate continual pretraining for another party, adding extra complexity.
To overcome this limitation, we propose a simplified and improved version of WSD, called WSD-S, for continual learning. Specifically, WSD-S resumes training immediately from the intermediate checkpoint using a high constant learning rate, avoiding the need to roll back to a previous checkpoint.
We evaluate the effectiveness of WSD-S with extensive experiments on LLMs from 0.1B to 1.2B parameters in a continual learning setting with 50B, 100B, and 200B tokens as the three target compute budgets. We empirically show that WSD-S has performance comparable with independent oracle runs with cosine learning rate schedules optimally tuned on each of the three budgets. Furthermore, WSD-S leads to a better validation loss than WSD under the same compute budgets due to the re-use of the decay period. We also show through ablation studies that the performance is relatively insensitive to the precise fraction of time spent decaying as long as it is near $10\%$ and the decay does not start shortly after a coincidental loss spike.
2 Related Work
Learning Rate Schedules. Learning rate schedules are crucial in deep learning, with previous studies exploring various options. Smith (2017) was the first to propose a cyclic triangular learning rate schedule that interleaves decreasing and increasing learning rates. Loshchilov & Hutter (2017) extended the idea to a cyclic cosine learning rate schedule. He et al. (2015) introduced the notion of warmup, which gradually increases the learning rate in the earlier training phase. Goyal et al. (2018); Hoffer et al. (2018); You et al. (2020) concluded that the learning rate should scale linearly with the batch size, which is further theoretically examined in Smith et al. (2020); Li et al. (2021); Malladi et al. (2023). You et al. (2019) performed an analysis on why learning rate schedules are helpful and suspected that the large learning rate at the beginning phase is mostly useful for avoiding memorization of noisy data, which is consistent with our analysis in Section 4.
In the LLM era, works including Hoffmann et al. (2022); DeepSeek-AI et al. (2024); Hu et al. (2024) examined how to choose learning rate schedules for pretraining. In particular, Hu et al. (2024) introduced a learning rate schedule called Warmup-Stable-Decay (WSD) that remains constant for the majority of the runs before decaying in language model pretraining, which were studied independently in Zhai et al. (2022); Ibrahim et al. (2024); Hägele et al. (2024).Raffel et al. (2023); Ibrahim et al. (2024) explored another possibility of using an inverse square root schedule to pretrain the language models. Rae et al. (2022); Gupta et al. (2023); Hu et al. (2024); Ibrahim et al. (2024) examined how to choose a learning rate schedule in a continual learning setting and verified that rewarming-up cosine learning rate brings performance drops that are costly to recover. A common belief is that the performance drop is due to the sudden increase in learning rate during rewarming-up. However, our work shows that increasing the learning rate after a short decay in WSD does not cause a similar performance drop as seen with the cosine learning rate, challenging the previous hypothesis. Instead, we suggest that the performance loss associated with rewarming-up cosine learning rate is due to the implicit damage it causes to the model, making it unsuitable for continual training. On the contrast, WSD avoids such damage by maintaining a high learning rate during the stable phase, hence the sudden increase in learning rate does not lead to performance drops in continual training.
Continual Learning. Continual learning, the process of updating the model with newly collected data, can improve the models’ knowledge and capability. Previous continual learning research (Aljundi et al., 2019; Veniat et al., 2021; Cossu et al., 2022; Dyer et al., 2022; Harun et al., 2023; Mehta et al., 2023) assumed significant domain shift and aimed to avoid forgetting old knowledge while learning new knowledge. Recent works including Hernandez et al. (2021); Lesort et al. (2023) suggested that optimizers including SGD and Adam have a knowledge accumulation effect and the effect of catastrophic forgetting may be less significant than expected, especially when replay is applied. Our work mainly focuses on continual pre-training without necessarily a strong domain shift and hence does not touch upon the effect of covariance shift. Continual learning is also extensively employed in large language models such as LLaMA to extend their capabilities, such as handling longer contexts (e.g., see Tworkowski et al. (2023); Peng et al. (2023); Chen et al. (2023); Dubey et al. (2024) and references therein) or dealing with new languages and domains (e.g., see Azerbayev et al. (2024); Rozière et al. (2024); Cui et al. (2024) and references therein).
Theoretical Understanding on Loss Landscape. A long line of research aims to better understand the loss landscape in deep learning (e.g., see Freeman & Bruna (2017); Garipov et al. (2018); Li et al. (2020) and references therein). We will highlight several phenomena that are related to our findings.
(1) Ill-conditioned directional sharpness and heavy-tailed noise: Zhang et al. (2020a; b) examined the gradient noise in language modeling and observed that the noise is heavy-tailed in multiple dimensions. Pan & Li (2023); Liu et al. (2024) showed that the loss has vastly different curvatures in different dimensions. Our river valley landscape is consistent with these findings.
(2) Benefit of large learning rates: Large learning rates have a provable regularizing effect in finding flatter minima (Kong & Tao, 2020; Wang et al., 2022), and flatter minima typically have a better generalization effect, even in the pretraining setting (Jiang et al., 2019; Blanc et al., 2020; Liu et al., 2022; Li et al., 2022; Ma et al., 2022; Lyu et al., 2023; Andriushchenko et al., 2023).
(3) Connecting loss landscape with feature learning: Some recent works (Nakkiran et al., 2019; Rosenfeld & Risteski, 2023) tried to understand how the loss landscape is formed through the lens of feature learning. Rosenfeld & Risteski (2023) showed that a large learning rate will cause oscillation in learning subtle classification rules while continuing to learn other more deterministic features. Wu et al. (2024); Cai et al. (2024) studied gradient descent dynamics on logistic regression, showing that a large learning rate will cause oscillation in the earlier phase but will lead to higher progress later in training. Wang et al. (2024) studied how to improve generalization and convergence by amplifying the update provided by the optimizer in the flat direction of the loss landscape. Pagliardini et al. (2024) developed a modification of the Adam optimizer based on optimization analysis on the Rosenbrock function, which is a special case of the river valley landscape.
The landscape analysis described in these previous works matches our river valley picture at a high level.
3 Theoretical Analysis with River Valley Loss Landscapes
Figure 3: A River Valley Loss Landscape and the Optimization Dynamics with Various Learning Rates.
Intuition and Outline. As illustrated in Figure 3, we will analyze the dynamics of different optimization algorithm on the river valley loss landscape. (1) The gradient descent (GD) iterates will quickly converge to the river and track the river afterward (Theorem 3.3). This holds for the continuous limit of GD, gradient flow, as well (Theorem 3.2). (2) Stochastic gradient descent (SGD) algorithm on a river valley loss landscape will bounce between the hillsides due to the noise instead of converging to the river, while still gradually progressing along the river over time (Theorem 3.4). A large learning rate results in significant oscillations but also facilitates rapid advancement in the river’s direction (the blue trajectory). Conversely, a small learning rate keeps the iterates closer to the river’s course but leads to slower movement along it (the yellow trajectory). While the oscillations in the mountain direction can cause a higher loss for a large learning rate in the short term, the large learning rate leads to faster fundamental progress along the river. (3) Finally, reducing the learning rate before termination (the cyan trajectory) helps mitigate these oscillations and reveal the underlying fundamental progress in the river’s direction (Theorem 3.5). In this scenario, the WSD algorithm is preferred: the stable phase allows for substantial progress along the river, and the final decay phase reduces oscillations, ensuring that the iterates terminate near the river’s course.
3.1 Setting and Assumptions
We will now formally present our theory. We use $w\in\mathbb{R}^{d}$ to denote the parameters and $L$ to denote the loss. Further, we use $\lambda_{k}\left(H\right)$ and $v_{k}\left(H\right)$ to denote the $k$-th largest eigenvalue and eigenvector of a matrix $H$, respectively. The “river” in the river valley is a 1-dimensional manifold $\mathcal{M}$ formalized below.
Assumption 1.
We assume the existence of a “river”, which is a 1-dimensional manifold $\mathcal{M}$ such that any point $w\in\mathcal{M}$ has a gradient $\nabla L(w)$ that is in the same direction as the minimal eigenvector direction of the Hessian, $v_{d}\left(\nabla^{2}L(w)\right)$.
Figure 4: Illustration of the Definition of the River.
Under this assumption, at every point on the river, the gradient $\nabla L(w)$ will align with the locally flattest direction, $v_{d}\left(\nabla^{2}L(w)\right)$, which we refer to as the river direction. All other directions orthogonal to the river direction are considered as the mountain directions, corresponding to the steep hillsides in our conceptual picture. This definition is illustrated in the 2-dimensional loss landscape shown in Figure 4. In this figure, the red line represents the top eigenvector (the sharpest direction), while the black line represents the smallest eigenvector (the flattest direction) at each point. Along the blue curve, which represents the river, the orange negative gradient aligns with the river direction. For a point on the hillside, the negative gradient pulls the point downward along the hillside towards the river, while also while also moving it along the river. This intuition will be rigorously formalized in the theorems that follow.
To analyze the optimization dynamics, we will consider a neighborhood $U$ of the river $\mathcal{M}$. Inside this neighborhood, we will impose several technical assumptions. For example, we will assume that the Hessian has an eigengap between the smallest and the second smallest eigenvalues, so that the river direction is well-defined in $U$. We will also assume that the last eigenvector of the Hessian $v_{d}$ changes slowly, which implies that the river direction does not change rapidly during optimization. We summarize these assumptions below.
Assumption 2 (Regularity Assumption).
There exists an open set $U$ containing $\mathcal{M}$ satisfying the following assumptions:
  1. 1.
    Analyticity. $L(w)$ is analytic with respect to $w$.
  2. 2.
    Bounded Hessian. There exists a constant $\gamma_{\mathrm{max}}>0$, such that $\forall w\in U,\|\nabla^{2}L(w)\|_{\textup{op}}\leq\gamma_{\mathrm{max}}.$
  3. 3.
    Existence of Eigengap. There exist constants $\gamma_{\mathrm{flat}},\gamma>0$, such that
    $\forall w\in U,\lambda_{d-1}\left(\nabla^{2}L(w)\right)>\gamma+4\gamma_{\mathrm{flat}},|\lambda_{d}\left(\nabla^{2}L(w)\right)|<\gamma_{\mathrm{flat}}.$
  4. 4.
    Slow Spinning of $v_{d}$. There exist constants $\Delta>0,\kappa\in[0,0.01)$, such that
    $\forall w\in U,\Delta_{\min}<\|\nabla L(w)\|_{2}\leq\Delta,\textup{ and }\|\nabla v_{d}\left(\nabla^{2}L(w)\right)\|_{\textup{op}}\leq\kappa\gamma/(2\Delta).$
    This means that the river direction $v_{d}$ changes slowly during optimization.
  5. 5.
    Uniqueness of $\mathcal{M}$. For any point $w\in U-\mathcal{M}$, the gradient $\nabla L(w)$ is not parallel to $v_{d}(\nabla^{2}L(w))$.
  6. 6.
    Conservation of Gradient Flows. There exists an open subset $V\subset U$ and a constant $r>\frac{10\Delta}{\gamma}$ for $\gamma$ defined in Assumption 2.3 such that $\forall w\in V$, the $r$-neighborhood of the gradient flow starting from $w$ stays in $U$ for continuous time $T_{\mathrm{max}}\geq{10\log(2\Delta/(\kappa\Delta_{\min}))}/{\gamma}$.
Throughout the analysis, $\kappa$ should be treated as a small dimensionless constant, indicating the river spins slowly. We can now define the river valley landscape.
Definition 3.1 (River Valley Landscape).
If a loss function $L$ satisfies Assumptions 1 and 2, then we will claim that the loss function is a river valley.
One simple example of a river valley landscape is the quadratic loss $L(x_{1},x_{2})=\frac{\gamma x_{1}^{2}}{2}-x_{2}$ with $\kappa$ equals to 00. In this case, the river is simply the line $x_{2}=0$. However, the river valley landscape can also be more complex and non-convex, see Figure 3 for an illustration. We will demonstrate that when the loss function is a river valley, the optimization trajectory of (stochastic) gradient descent will track the river. In fact, we can even prove that the iterates will follow the river with a predictable pace, which is characterized by the reference flow defined below.
Reference Flow. We introduce a Riemannian gradient flow constrained to the river $\mathcal{M}$, serving as a reference in the following theorems. This flow intuitively represents the dynamics of iterates during a gradient flow on the loss constrained by the river. We will denote the projection to the tangential space of the river as ${P}_{\mathcal{M}}(w)$ for $w\in\mathcal{M}$ and choose an arbitrary starting point $x_{0}$ on the river. The reference flow is defined as
$dx(T)=-{P}_{\mathcal{M}}\left(x\left(T\right)\right)\nabla L(x(T))dT,\quad x(0)=x_{0}.$ (1)
Here, we use $x$ to represent a point on the river, distinguishing it from $w$, which denotes a weight in the original space. $T$ refers to the continuous time variable.
3.2 Gradient Flow Dynamics
In this section, we will consider gradient flow in the river valley landscape starting from a point $w\in V$, with $V$ defined in Assumption 2.6:
$dw(T)=-\nabla L(w(T))dT,\quad w(0)=w\in V.$ (2)
We prove in the theorem below that the gradient flow starting from $w$ will eventually converge near the river and remain close to it. Subsequently, if we project the iterate $w(T)$ onto the river, the projection will move along the river at a pace similar to the reference flow $x(\cdot)$ (eq. 1). This phenomenon is visualized on Figure 5(a).
Theorem 3.2.
If a loss $L$ is a river valley (Definition 3.1), for the gradient flow $w(T)$ defined in Equation 2, the iterate will obey the following dynamics:
  1. 1.
    The iterate will first converge to a neighborhood of the river. The distance betweeen the iterate and the river is bounded after a constant converge time $T_{\mathrm{converge}}={2\log(2\Delta/(\kappa\Delta_{\min}))}/{\gamma}$.
    $\mathrm{dist}(w(T_{\mathrm{converge}}),\mathcal{M})=\mathrm{min}_{T}\|x(T)-w(T_{\mathrm{converge}})\|_{2}\leq 2\kappa\Delta/\gamma.$
  2. 2.
    After convergence, the iterate will track the river closely with the same pace as the reference flow. There exists a time shift $T_{0}$ depending on $w(T_{\mathrm{converge}})$, such that for any $T\in[T_{\mathrm{converge}},T_{\mathrm{max}}]$ for $T_{\mathrm{max}}$ defined in Assumption 2.6, there exists a $\tilde{T}\in[(1-\epsilon)T,(1+\epsilon)T]$ satisfying that,
    $\|x(T_{0}+\tilde{T})-w(T)\|_{2}$ $\leq 2\kappa\Delta/\gamma,$
    for $\epsilon=30\kappa$.
(a) Gradient Flow Dynamics
(b) Gradient Descent Dynamics
(c) Stochastic Gradient Descent
Figure 5: Validation of Theory on a 2D Function. We validate Theorems 3.2, 3.3 and 3.4 using a 2D example with the loss function $L(x_{1},x_{2})=(x_{2}-\sin(x_{1}))^{2}+0.2|10-x_{1}|$. The blue curve represents the "river", where the gradient aligns with the minimal eigenvector of the Hessian. (1) On the left, we observe that multiple randomly initialized gradient flows converge near the river and follow it closely thereafter, consistent with Theorem 3.2. (2) In the middle, we show that discrete gradient descent with a learning rate $\eta=0.6$ shows similar behavior: after initial oscillations, the gradient descent iterates align closely with their projections on the river. The inset illustrates that the $t$-th projection’s progress along the reference flow (eq. 1) approximately equals $\eta t$, as predicted by Theorem 3.3. (3) On the right, we further illustrate that stochastic gradient descent (SGD) also tracks the river. In contrast to the discrete-step gradient descent, the iterates oscillate around the river rather than staying on it. The trajectory with a larger learning rate exhibits both faster progress and greater oscillations compared to the trajectory with a smaller learning rate, as predicted by Theorem 3.4.
The proof is deferred to Section A.4. In this theorem, the lower bound on $T$ represents the time required for the iterate to converge near the river. Here $x(T_{0}+\tilde{T})$ can be viewed as a projection of $w(T)$ onto the river. As both the geometric error ($2\kappa\Delta/\gamma$) and the time-alignment error ($\epsilon$) vanish when $\kappa$ is small, this projection is not only close to $w(T)$ but also moves at nearly the same rate as the reference flow. Here the term $T_{0}$ acts as a shift, reflecting the dependency on the initialization, as optimization trajectories starting from different initial points will enter the river at distinct locations. The term $\tilde{T}$ represents the progress made along the river. This interpretation will remain consistent in the subsequent sections.
Proof Sketch. The proof of Theorem 3.2 relies heavily on the structure of the Hessian matrix. The eigengap assumption (Assumption 2.3) ensures that the loss constrained to the mountain dimensions is steep and convex enough to guarantee convergence to the river. The slow spinning assumption (Assumption 2.4) ensures that the river direction does not change rapidly, allowing the iterate to follow the river closely after convergence.
3.3 Gradient Descent Dynamics
We will now analyze gradient descent with a discrete learning rate. This process is inherently discrete, differentiating it from the continuous gradient flow analyzed above. Similar to the continuous case, an iterate far from the river will converge to the river (as visualized in the first few steps of Figure 5(b)). To ease our analysis, we will skip the convergence analysis and assume the starting point $w$ lies on the course of the river.
$w_{\eta}(t+1)-w_{\eta}(t)=-\eta\nabla L(w_{\eta}(t)),\quad w_{\eta}(0)=w\in\mathcal{M}.$ (3)
Here we use $t$ to denote the discrete time step, in contrast to the continuous time variable $T$ used in the previous section. When the learning rate $\eta$ is small, despite the discretization error, gradient descent will still track the river closely at approximately the same pace as the reference flow similar to gradient flow. In this case, the progress over $t$ steps will be approximately $\eta t$, as shown in the following theorem.
Theorem 3.3.
If a loss $L$ is a river valley (Definition 3.1), when $\eta<\frac{\gamma}{2\gamma_{\mathrm{max}}^{2}}$, for the gradient descent $w_{\eta}(T)$ defined in Equation 2 with initialization $w$ on the river, there exists a time shift $T_{0}$ depending on $w$ and $\eta$, satisfying that for any $\eta t\leq T_{\mathrm{max}}$, there exists a $\tilde{T}\in[(1-\epsilon)\eta t,(1+\epsilon)\eta t]$ satisfying that,
$\|x(T_{0}+\tilde{T})-w_{\eta}(t)\|_{2}$ $\leq 10\kappa\Delta/\gamma\,,$
for $\epsilon=30\kappa+4\eta\gamma_{\mathrm{flat}}$.
The proof is deferred to Section A.5. We observe that the distance of the iterates from the river remains on the same order as in Theorem 3.2. However, the learning rate will affect the pace of the iterates, introducing the additional error term $\eta\gamma_{\mathrm{flat}}$ in time-alignment error ($\epsilon$). This term is unavoidable, as the discretization inherent in gradient descent introduces a slight deviation from the continuous gradient flow, even in the absence of the hill component of the loss. Finally, the Theorem 3.3 predicts that a larger learning rate $\eta$ will induce higher progress $\eta t$ down the river given the same number of steps $t$.
We illustrate Theorem 3.3 in Figure 5(b). Here the red points correspond to gradient descent iterates and green points are the corresponding predicted projected points on the river. As the theory predicts, the distance between the iterates and projections diminishes after a few steps. Finally, the inset indicates that the projected points traverse the reference flow on the river at a rate proportional to the learning rate $\eta$.
3.4 Stochastic Gradient Descent Dynamics
The above analysis holds for deterministic dynamics and we will now proceed to model the stochasticity in the optimization process. This stochasticity will stop the iterate from fully converging to the river and lead to oscillation in the mountain direction. To simplify the analysis, we will consider a special case where the river direction is a constant and the river reduces to a straight line. We conjecture that the results below can be generalized to a more general setup and leave the detailed analysis to future work.
Assumption 3 (Straight River).
For $U$ in Assumption 2, $\forall w\in U,\|\nabla v_{d}(\nabla^{2}L(w))\|_{2}=0$. In this case, the river is a straight line parallel to the direction of $v_{d}(\nabla^{2}L(w))$.
Under Assumption 3, $v_{d}(\nabla^{2}L(w))$ is a constant vector for $w\in U$ and we will use $v_{d}$ to denote this vector. We will also assume that the update is deterministic in the direction of the river, which simplifies our proof while still capturing the essential dynamics of SGD. Consequently, we can express the SGD update as follows:
$\tilde{w}(t+1)=\tilde{w}(t)-\eta_{t}\nabla L\left(\tilde{w}\left(t\right)\right)+\eta_{t}{g}_{t},\quad{g}_{t}\sim\mathcal{N}\left(0,\sigma^{2}\left(\mathcal{I}_{d}-v_{d}v_{d}^{T}\right)\right),\quad\tilde{w}(0)=w\in\mathcal{M}.$ (4)
Here $\mathcal{N}\left(\mu,\Sigma\right)$ indicates the normal distribution with mean $\mu$ and covariance $\Sigma$. Compared to deterministic gradient descent, the introduced noise ${g}_{t}$ causes the iterates to deviate from the river instead of fully converging to it (see the difference between Figure 5(b) and Figure 5(c)). Consequently, we need to impose additional assumptions on the loss landscape to analyze the dynamics of SGD.
Assumption 4 (Regularity Assumption for SGD).
In the setting of Assumptions 2, we assume in addition the following:
  1. 1.
    Bounded Hessian. There exists a constant $\tau>0$, such that for any weight $w\in U$, the nuclear norm of the Hessian is bounded.
    $\|\nabla^{2}L(w)\|_{*}=\sum\limits_{i=1}^{d}|\lambda_{i}\left(\nabla^{2}L(w)\right)|\leq\tau.$
  2. 2.
    Bounded Third Order Gradient. There exist constants $\rho>0,\kappa^{\prime}\in[0,0.01]$, such that,
    $\|\nabla^{3}L(w)\|_{\textup{op}}\leq\rho,\quad\Delta\rho\leq\kappa^{\prime}\gamma^{2}.$
  3. 3.
    Bounded Loss. There exists a constant $M>0$ such that $\forall w,L(w)<M$.
In this assumption, we treat $\kappa^{\prime}$ as a small constant, indicating that the influence of the third-order gradient is minimal. This suggests that the overall shape of the loss landscape is predominantly governed by the first and second-order information. We will now analyze the dynamics of SGD in two phases: the stable phase and the decay phase. In both phases, the iterates will track the reference flow with progress linear to the sum of the learning rates. At the same time, the iterates will also incur a loss due to the noise, which will be linear with respect to the learning rate on the last step.
3.4.1 Stable Phase
We start with the stable phase, where the learning rate $\eta_{t}=\eta$ remains constant. Similar to the deterministic case, we will demonstrate that the expected loss of the iterate $\mathbb{E}[L(\tilde{w}(t))]$ closely follows the loss along the reference flow $L(x(T))$, with $T\approx\eta t$. However, the stochasticity in the updates introduces an additional term, $\eta\sigma^{2}$, into the loss. This provides a formal basis for decomposing the loss into its river and hill components, that is, $L(x(T))$ and $\eta\sigma^{2}$, respectively.
Theorem 3.4.
Suppose a loss $L$ is a river valley (Definition 3.1) and satisfies Assumptions 3 and 4. Then, for any constants $\delta\in(0,1)$ and $T\leq T_{\mathrm{max}}$, for sufficiently small learning rate $\eta$ depending on the regularity constants,  [^1] Deferred to Assumption 7 in Appendix. the SGD iterates (defined in Equation 4) with $\eta_{t}=\eta$ satisfies that for any integer $t\in[1/{\eta\gamma},T/\eta]$, there exists a $\tilde{T}\in[(1-\epsilon_{t})\eta t,(1+\epsilon_{t})\eta t]$ satisfying that,
$\mathbb{E}[L\left(\tilde{w}(t)\right)]-L(x(\tilde{T}))$ $={(d-1)\eta\sigma^{2}}/{2}+\epsilon_{L}$
where $\epsilon_{t}=4\eta\gamma_{\mathrm{flat}}$ and $|\epsilon_{L}|\leq\tau\eta^{2}\sigma^{2}+\rho(C{d\eta\sigma^{2}}/{\gamma})^{3/2}+C\kappa^{\prime}d\eta\sigma^{2}+\delta(2M+\eta\sigma^{2}d)\ll(d-1)\eta\sigma^{2}$ with $C=200\log(64\gamma T/\delta)$.
The proof is deferred to Section A.6. In Theorem 3.4, the error term in the approximation of the pace of the projection remains the same as in the deterministic case (Theorem 3.3). However, the stochasticity introduces an additional hill component $(d-1)\eta\sigma^{2}/2$ to the expected loss at the iterate. The hill component increases linearly with the learning rate. The error term in loss term $\epsilon_{L}$ can be decomposed into three parts: (1) $\tau\eta^{2}\sigma^{2}+\rho(C{\eta\sigma^{2}}/{\gamma})^{3/2}$ are higher order discretization effects of learning rate $\eta$; (2) $C\kappa^{\prime}d\eta\sigma^{2}$ is caused by the change of the Hessian in the valley dimensions and will diminish when $\kappa^{\prime}$ is small; (3) $\delta(2M+\eta\sigma^{2}d)$ accounts for the small chances that the iterate will escape the neighborhood of the river due to the stochastic updates. While the theorem only considers the case where $v_{d}$ is a constant vector, we conjecture that the theorem can be extended to a general setting and verify this conjecture on a toy loss (see Figure 5(c)).
3.4.2 Decay Phase
Finally, we will consider the decay phase in training and will show that a proper decaying schedule can reduce the hill component of the loss rapidly. We will first define our decaying schedule, starting from step $t_{s}=\left\lceil T/\eta\right\rceil$:
$\eta_{t}=\frac{\eta}{2+(t-t_{s})\eta\gamma},\quad$ $t_{s}\leq t\leq 1.1t_{s}.$ (5)
We choose this schedule to maximize the loss decrease rate on a quadratic function (see Section A.2) because we perform quadratic approximations of the loss near the river in our analysis. Our theorem predicts that the hill component of the loss will decrease linearly with the learning rate under this learning rate schedule, consistent with the empirical findings in Hu et al. (2024).
Theorem 3.5.
Under the setting of Theorem 3.4, the SGD iterates (defined in Equation 4) with the learning rate schedule defined in Equation 5 satisfies that for any integer $t\in[t_{s},1.1t_{s}]$, there exists a $\tilde{T}\in[(1-\epsilon_{t})T(t),(1+\epsilon_{t})T(t)]$ satisfying that,
$\mathbb{E}[L\left(\tilde{w}(t)\right)]-L(x(\tilde{T}))$ $\leq{(d-1){\color[rgb]{1,0,0}{\eta_{t}}}\sigma^{2}}/{2}+\epsilon_{L}$
with $T(t)=T+\sum\limits_{k=t_{s}}^{t}\eta_{k}$.
Figure 6: Predicted Loss Curve of SGD By Theorems 3.4 and 3.5 on Loss $L(x_{1},x_{2})=\gamma x_{2}^{2}/2-x_{1}$.
The formal proof is deferred to Section A.7. Compared with Theorem A.30, the hill component is now dominated by $(d-1)\eta_{t}\sigma^{2}/2$, scaling linearly with the decaying learning rate. When the oscillation level $\sigma$ is large compared to the loss changes along the river, the loss decrease can then appear faster in the decay phase than in stable phases (see Figure 6). Further, the decaying phase also makes progress along the river, which corresponds to the term $\sum\limits_{k=t_{s}}^{t}\eta_{t}$ in the theorem. Finally, the terms used in our theorem match the scaling law formulation in the concurrent work (Tissue et al., 2024).
3.5 Future Extensions
We only consider a 1-dimensional river in the theoretical analysis above. However, it is possible to extend our assumptions to a generalized $k$-dimensional river, where the gradient lies in the flattest $k$-dimensional subspace for every point on the river.
Assumption 5.
We assume the existence of a “generalized river”, which is a $k$-dimensional manifold $\mathcal{M}$ such that any point $w\in\mathcal{M}$ has a gradient $\nabla L(w)$ lies in the eigenspace spanned by the last $k$ eigenvectors’ direction of the Hessian, $\{v_{i}\left(\nabla^{2}L(w)\right)\mid i\in[d-k+1,d]\}$.
Based on this assumption, we can define generalized river valley Structure. We believe that under some appropriate assumptions, there will be a similar coupling between the dynamics in the original river valley landscape and the dynamics constrained on the generalized river, and leave this for future work.
There are several other potential technical improvements to our theory that could be explored in future works. First, the analysis of the stochastic setting may be extended to include a river that is not a straight line. Second, the multiplicative constant of 4 in the eigengap condition (Assumption 2.3) might be eliminated through more refined analysis. Lastly, the upper bound on the learning rate in Theorem 3.3 (currently $\gamma/2\gamma_{\mathrm{max}}^{2}$) could potentially be improved to $\Theta(1/\gamma_{\mathrm{max}})$ through more fine-grained analysis, which is the maximal learning rate in the quadratic case.
3.6 Visualizing the River Valley
We use a direct probing method to verify our theory. Our theory suggests that when the learning rate is large, the model will bounce back and forth between the sharp valleys. However, in the decay phase, the model will move downwards the hillside to approach the river. This suggests that if we connect two checkpoints in the stable phase, we should expect to see a projection of the valley, and if we connect two checkpoints in the decay phase, we should expect to see smooth decreasing curves. To verify this, we pretrain a 124M GPT-2 model on OpenWebText. In the first run, we train the model with a constant learning rate for 25B tokens and interpolate between two checkpoints at 20B and 25B tokens (Figure 7(a)). In the second run, we branch off from the first run at 20B tokens and decay the learning rate for 5B tokens, and we interpolate between two checkpoints at 20B and 25B tokens (Figure 7(b)). The interpolation results closely resembles our theory. This observation is also consistent with Sanyal et al. (2023) which shows weight averaging improves model performance in the earlier part of the cosine training runs, where the learning rates are higher. Additionally, the smooth decreasing curves we observed when connecting two checkpoints in the decay phase are consistent with the findings in Hägele et al. (2024).
(a) Stable Phase
(b) Decay Phase
Figure 7: Probing Loss Landscape. We validate the river valley analogy by interpolating stable and decay phases in GPT-2 pretraining experiments. We observe that loss resembles a valley when constrained on the segment connecting two models during the stable phase and smoothly decreases when connecting two models during the decay phase.
4 Uncertainty Variation in Data Distribution Shapes the River Valley Landscape
What causes the loss landscape to resemble a river valley structure? In this section, we propose and validate the hypothesis that variations in next-token uncertainty shape the loss landscape. When predicting a deterministic fact, a large learning rate can boost the model’s confidence, accelerating learning. However, when the next token is inherently ambiguous—such as the continuation of a phrase like "I am"—the model must learn a calibrated distribution, which may necessitate a smaller step size. This variation in uncertainty leads to differences in sharpness across the loss landscape, resulting in the river valley structure.
A Toy Bigram Language. We formalize this intuition using a synthetic language composed of cities and names, where each city corresponds to a unique distribution of its citizens’ names. For instance, one city might have a highly deterministic distribution, with most residents named "Ken“, while another city may have a more diverse distribution of names. This synthetic language follows the structure in Allen-Zhu & Li (2024). The goal is to learn the distribution of names conditioned on each city. We show that cities with more deterministic name distributions align with flatter regions in the loss landscape (the "river"). In contrast, cities with more diverse name distributions correspond to sharper regions (the "hillsides").
Figure 8: Visualization of Toy Bigram Language. We design a synthetic dataset where each city has a unique name distribution. The left shows the name distributions for two cities, one deterministic and one stochastic.
Formally, let the set of cities be represented by $\{1,\ldots,n\}$ and the set of names by $\{1,\ldots,m\}$. Data is generated by first selecting a city $i$ uniformly at random, then sampling a name $j$ according to the city’s name distribution. The name distribution for city $i$ is parameterized by a categorical distribution $\mathrm{Categorical}([P_{i,j}]_{j=1}^{m})$, where $P_{i,j}$ represents the probability of selecting name $j$ in city $i$, and each $P_{i,j}>0$. To quantify the uncertainty in each city’s name distribution, we compute the Gini impurity of the distribution as:
$U_{i}=\mathrm{I_{G}}(\mathrm{name}\mid\mathrm{city}=i)=1-\sum_{j=1}^{m}P_{i,j}^{2}\in\left[0,1-\frac{1}{m}\right).$
The value of $U_{i}$ reflects the uncertainty of city $i$’s name distribution. When the distribution is close to deterministic—i.e., there exists a $j$ such that $P_{i,j}$ is near 1—$U_{i}$ approaches its lower bound of 0. Conversely, for a nearly uniform distribution, $U_{i}$ approaches its upper bound of $1-\frac{1}{m}$. Given this setup, we parameterize our model with $\Theta\in\mathbb{R}^{n\times m}$, where each row corresponds to a city and each column to a name. The model estimates the probability of name $j$ for city $i$ using the softmax function $\frac{\exp(\Theta_{i,j})}{\sum_{k=1}^{m}\exp(\Theta_{i,k})}$.
We use sampled data to train this model. Specifically, for each sampled city $i$, a name $j$ is drawn according to the true distribution $\mathcal{P}_{i,j}$. The model’s task is to predict the correct name based on the sampled city, minimizing the negative log-likelihood (NLL) over the sampled data. The population loss is given by:
$L(\Theta)=\frac{1}{n}\sum_{i=1}^{n}\ell_{i}(\Theta_{i,:}),\quad\ell_{i}(\Theta_{i,:})=-\sum_{j=1}^{m}\mathcal{P}_{i,j}\log\frac{\exp(\Theta_{i,j})}{\sum_{k=1}^{m}\exp(\Theta_{i,k})}.$ (6)
This loss is separable across different cities, meaning that the contribution from each city is independent. The loss component $\ell_{i}(\Theta_{i,:})$ captures the contribution from city $i$, and different name distributions across cities lead to different forms of $\ell_{i}$. Considering a parameter $\Theta^{*}$ that minimizes the loss $L$, we will show that cities with more stochastic name distributions correspond to sharper components in the loss landscape, as reflected by the average-direction sharpness of $\ell_{i}$ ($\mathrm{Tr}(\nabla^{2}\ell_{i}(\theta))\mid_{\theta=\Theta^{*}_{i,:}}$).
Lemma 4.1.
The average-direction sharpness of loss component $\ell_{i}$ at $\Theta^{*}$ equals the uncertainty of the name distribution ($U_{i}$). $\mathrm{Tr}(\nabla^{2}\ell_{i}(\theta))\mid_{\theta=\Theta^{*}_{i,:}}=U_{i}.$
Lemma 4.1 demonstrates that at the global minimum, the sharpness associated with a city decreases as the city’s name distribution becomes more deterministic. This aligns with the intuition that a deterministic token corresponds to a flatter loss direction.
In a generalized river valley landscape, the hillsides represent the sharper components of the loss, which in this context correspond to cities with more stochastic name distributions. Conversely, the river corresponds to cities with more deterministic name distributions. We can further establish the existence of a generalized river (Assumption 5) in this loss landscape under appropriate assumptions about $\mathcal{P}$ (see Theorem A.33). Along the river, the gradient remains nonzero only for the cities with more deterministic name distributions, reinforcing the connection between determinism and flatness in the loss landscape.
Figure 9: Loss Curves and Relative Loss Decrease Distribution. The left figure reproduces the non-traditional loss curve of WSD on the synthetic language, showing that the loss remains elevated during the stable phase but sharply declines during the decay phase. The right figure displays the distribution of the relative loss difference from the two different distributions comparing the stable phase and the decay phase. We can observe that the loss decrease over the stochastic distribution is much larger than the decrease in the deterministic distribution.
Empirical Verification. We empirically verify that the loss curve of WSD can be reproduced in our synthetic setting. The dataset used contains two types of cities: (1) a deterministic type with name distribution’s entropy less than 0.2, and (2) a stochastic type with name distribution’s entropy greater than 1. Each type contains 1.8k cities and there are 10 possible names. We first train a toy model on this synthetic data and successfully replicate the non-traditional loss curve of WSD (Figure 9,left).
Next, we convert the data into a synthetic language in the format "The resident of [CITY]: [NAME]" and fine-tune a GPT-2 model, pretrained on OpenWebText, using this synthetic data. We experiment with two different learning rate schedules: a constant schedule (stable) and a decaying schedule (decay). We then calculate the difference in loss between the two models’ predictions for the first token of "[NAME]". A significant Spearman correlation of 0.388 (Figure 9,right) is observed between the loss difference and the ground truth entropy per city. This correlation indicates that the loss decrease is greater during the decay phase for more stochastic populations. Furthermore, although the decay phase achieves a lower overall loss, the mean loss for the deterministic sub-population is slightly higher than in the stable run, suggesting that the stable run better learns the deterministic sub-population.
5 WSD-S: A Simplification of the WSD Schedule
Figure 10: Comparison with the Cosine Oracles. We show that the WSD-S schedule can perform similarly to the Cosine schedules in a single run. The $\star$ in the graphs visualize the terminating validation loss of different Cosine runs. The largest validation loss gap between the WSD-S and the Cosine schedules is 6e-3. The lower right figure plots the learning rate curves used in this experiment.
5.1 Method
The goal of continual pretraining is to generate checkpoints that exhibit good performance at multiple compute budgets in one run. Formally, our goal is to achieve multiple intermediate checkpoints $\theta_{T_{k}}$, each corresponding to a computing budget (number of steps) $T_{k}$ for $k\in\{1,\ldots,K\}$.
A strong baseline to measure the performance of $\theta_{T_{k}}$ would be running cosine learning rates (Figure 10, lower right) for each budget $T_{k}$ separately, decays the learning rate linearly to the cosine function between $[0,\pi]$. We will dub this oracle method as Cosine-Oracle. However, Cosine-Oracle can’t be done in a single run and will incur a high total compute budgets $\sum_{k}T_{k}$. A simple modification to Cosine-Oracle is to use multiple consecutive cosine learning rates between $T_{k-1}$ and $T_{k}$ (Figure 2(b), last row), which we will dub as Cyclic-Cosine. Cyclic-Cosine only requires a total compute budgets $T_{K}$. However, it leads to non-negligible performance lost compared to Cosine-Oracle (Hu et al. (2024)).
Warmup-Stable-Decay (WSD) address this issue by maintaining a main branch that keep using a constant learning rate after warmup process and branch off using a decaying learning rate to achieve intermediate checkpoints. One can then continue pretrain from a checkpoint in the main branch by resuming with the same constant learning rate. Formally, WSD introduce decay starting points $D_{1},\ldots,D_{k}$ such that $T_{i-1}<D_{i}<T_{i}$. WSD will then correspond to the following process (Figure 2(b), second row):
  1. 1.
    Get a main branch of checkpoints ${\theta^{\mathrm{main}}}$ by running a constant learning rate schedule for $D_{K}$ steps.
  2. 2.
    For each $k$, run a decaying learning rate schedule for $T_{k}-D_{k}$ steps starting from $\theta^{\mathrm{main}}_{D_{k}}$ to get $\theta_{T_{k}}$.
The above process reutilizes the main branch of checkpoints ${\theta^{\mathrm{main}}}$ for each $T_{k}$ and hence reduces the total compute budget to $T_{K}+\sum_{k}(T_{k}-D_{k})$. Recall that in the river valley landscape model, the Warmup-Stable-Decay (WSD) algorithm can be viewed as a combination of a large learning rate phase to speed up progress down the river and a rapid learning rate drop at the end to reduce the oscillation. Because the decay phase also makes progress along the river (see Theorem 3.5), we proprose a simplified version of WSD, called Warmup-Stable-Decay-Simplified (WSD-S), that continues with another stable phase leaving off of the end of the previous decay phase (see the first row of LABEL:{fig:learningrate}) without separating the training process into two branches. Formally, the WSD-S learning rate schedule is defined as follows:
$\eta_{t}=\begin{cases}\text{decay}(T_{i}-D_{i},\eta_{\mathrm{max}},\eta_{\mathrm{min}})[t-D_{i}]&\text{if }\exists i,D_{i}<t\leq T_{i};\\ \eta_{\mathrm{max}}&\text{otherwise.}\\ \end{cases}$ (7)
The key difference from our methods is the choice of initialization point when retraining starts. In WSD, the second stable phase uses the model before the decay phase, whereas we use the model after it. This process is more convenient to implement because it does not require rolling back to the main branch after each decay phase. Here the learning rate decay function $\mathrm{decay}$ can take many forms that decay the learning rate from $\eta_{\mathrm{max}}$ to $\eta_{\mathrm{min}}$ over $T_{i}-D_{i}$ steps. In this paper, we will use the inverse proportional decay function as defined in Equation 8 for all experiments (visualized in Figure 2(b), first two rows).
$\frac{1}{\text{decay}(T,\eta_{\mathrm{min}},\eta_{\mathrm{max}})}$ $=\left[\frac{t}{T}\frac{1}{\eta_{\mathrm{min}}}+\left(1-\frac{t}{T}\right)\frac{1}{\eta_{\mathrm{max}}}\ \Bigg{|}\ t\in\{0,1,\ldots,T\}\right].$ (8)
This function is motivated by the analysis on quadratic functions in Theorem 3.5. The reciprocal of the learning rate linearly interpolates from the reciprocal of maximal to the reciprocal of minimal learning rate.
5.2 Experiments
5.2.1 Setup
Architecture and data. We adopt the LLaMA architecture from Touvron et al. (2023), adjusting the hyperparameters to create four model sizes: 0.1B, 0.3B, 0.6B, and 1.2B. The exact hyperparameters are deferred to Appendix B. These models are trained on the Pile dataset (Gao et al., 2020) with a context length of 4096 and a batch size of 4M tokens.
Figure 11: Cosine Learning Rate Implicitly Hurts the Models for Future Continual Learning. We show that while WSD and the cosine learning rate schedule may produce similar validation loss in a single run, a model trained with the cosine learning rate schedule is implicitly hurt compared to the model trained with WSD for future Continual learning. On the 0.6B models, after training the models for 50B tokens using both WSD-S and the cosine learning rate schedule, we continually train two models for another 50B tokens using both learning rates. We observe that the model trained with WSD-S consistently outperforms the model trained with the cosine learning rate when used as the starting point for further training.
Implementation. We set the batch size to 1024 and fixed the peak learning rate for the same model size for all the methods. For the 0.1B and 0.3B models, we use a peak learning rate of 6e-4, and for the 0.6B and 1.2B models, we use a peak learning rate of 4e-4. These values are chosen following current empirical practice (e.g. see Groeneveld et al. (2024)). We set the minimal learning rate to 0.1 of the peak learning rate. We use a TPU v3-256 model to train the model with the Levanter framework in Jax (Bradbury et al., 2018; CRFM, 2024). The fraction of time spent decaying is chosen to be $10\%$. The only exception is that when running WSD on the 0.3B models, we encounter a loss spike after training for 22.5B tokens and decay at the checkpoint trained for 22B tokens instead. This change is in favor of WSD in our comparison between WSD-S and WSD. The detailed hyperparameters are deferred to Appendix B.
5.2.2 Results
Figure 12: Comparison With WSD. We show that WSD-S performs favorably compared with WSD when the total computes is fixed, achieving a consistent improvement over WSD on all the model sizes when trained for approximately 200B tokens.
WSD-S performs competitively with Cosine-Oracle. The three endpoints of WSD-S are set at 50B, 100B, and 200B tokens for all models. As shown in Figure 10, WSD-S delivers competitive results compared to Cosine-Oracle in a single run.
WSD-S significantly outperforms Cyclic-Cosine. We compare the Cyclic-Cosine and the WSD-S on 0.6B models with a total token budget of 100B tokens. Both schedules reduce to a minimal learning rate at 50B tokens to obtain an intermediate checkpoint. Our results show that WSD-S outperforms Cyclic-Cosine with a significant performance gap of 4e-2 (Figure 11). A common belief is that loss spiking after increasing the learning rate is the main cause of the performance loss in Cyclic-Cosine. However, this belief does not explain the advantage of WSD-S. We hypothesize that a model trained with a small learning rate for too long, as with Cosine, is implicitly hurt compared to a model trained with a large learning rate for the majority of the run, as with WSD or WSD-S.
To show that the model trained with WSD is more suitable for continual training, we conducted ablation studies by interchanging the schedules in the latter half of the runs to create two new learning rates (Cosine-WSD and WSD-Cosine). Among the four runs, the model trained using WSD for the first half consistently achieved lower loss in continual learning, indicating that WSD produces models more suitable for continual learning, even after learning rate decay.
[…]
(a) 1.2B Models on 200B tokens
(b) 0.1B Models on 100B tokens
(c) 0.1B Models on 100B Tokens. Decay Near a Loss Spike (6%)
Figure 13: Ablation Study on the Sensitivity of Fraction of Time Decaying. This study examines two settings: a smaller scale with 0.1B parameters trained on 100B tokens (middle figure) and a larger scale with 0.6B parameters trained on 200B tokens (left figure). The results indicate that the final performance is similar when the decay phase is 8%-12% of the total training steps. However, the right figure demonstrates a significant performance loss when decaying near a loss spike. It compares two training loss curves with decay phases of 8% and 6% of the total compute on the 0.lB models, where the latter starts immediately after a loss spike, leading to a validation loss increase of 2e-2.
WSD and WSD-S is not sensitive to the fraction of time spent decaying We conclude with an ablation study on the fraction of time spent decaying, and the result is shown in Figure 13. The final performance matches tightly within the range of $8\%$ to $12\%$, showing small sensitivity to the choice of the decay portion.
However, in our experiments, we observe that decaying near a loss spike can lead to a significant performance loss (Figure 13,right).With the large learning rate, the training runs tend to be very volatile and there are multiple loss spikes in the training (see Figure 10). If a decay happens closely after a loss spike and the loss has not yet decreased to its original level, it is typical that the final validation loss will be worse by 1e-2 or even more. We observe the same phenomenon for WSD, and when such a scenario happens, we suggest either running longer till the loss stabilizes or rolling back to a slightly earlier checkpoint before the loss spikes and decays from there.
6 Conclusion
In this work, we propose a river valley landscape metaphor to explain why large learning rates can make implicit progress in training that can only be revealed when the learning rate is decayed. Further, we propose a simple data model that can generate the river valley landscape, attributing the difference of sharpness in different directions to the difference in uncertainty between different tokens. Based on our theory, we propose a learning rate schedule called WSD-S that can produce multiple intermediate checkpoints in a single-consecutive run as good as if we prespecified the individual budgets and used a tuned cosine learning rate. The learning rate schedule involves rapidly decaying to a minimal learning rate level at these budgets and resuming to a high constant learning rate immediately afterward and keeping the learning rate constant in the rest of the runs. From training language models from 0.1B to 1.2B parameters, we show WSD-S performs competitively with cosine learning rate and outperforms other compute-agnostic learning rates including WSD and cyclic-cosine learning rate fixing the total compute budget.
Appendix A Omitted Proofs
A.1 Notation.
To denote $a^{T}b$ for two vectors, we will $\langle a,b\rangle$. We will use the following function to denote the directional derivative of a mapping $F:\mathbb{R}^{d}\to\mathbb{R}^{m}$:
$\nabla F(x)[v]=\lim_{\alpha\to 0}\frac{F(x+\alpha v)-F(x)}{\alpha}.$
A.2 A warmup on the quadratic function
We will first motivate the decaying function we choose Equation 8 using a simple example on quadratic function.
Lemma A.1.
Assuming that we are considering the following gradient descent
$y_{t+1}=y_{t}-\eta_{t}\nabla(\gamma y_{t}^{2}/2)-\eta{g}_{t},{g}_{t}\in\mathcal{N}(0,\sigma^{2}\mathcal{I}).$
Suppose $\eta_{0}=\eta_{\mathrm{max}}$ and $y_{0}$ follows a normal distribution $\mathcal{N}(0,\eta_{\mathrm{max}}\frac{\sigma^{2}}{2\gamma-\eta_{\mathrm{max}}\gamma^{2}})$. Then the following two statements hold,
  1. 1.
    If $\forall t,\eta_{t}=\eta_{0}$, $y_{t}$ will follow the same distribution as $y_{0}$.
  2. 2.
    Consider all the learning rate schedule $\eta_{t}$, the following is the optimal
    $\forall t\geq 1,\eta_{t}^{*}=\frac{1}{\gamma(t-1)+\frac{2}{\eta_{\mathrm{max}}}}$
    in the sense that it yields the fastest expected loss decrease. Suppose $\eta_{t}^{*}$ corresponds to iterates variables $y_{t}^{*}$, for any $\eta_{t}$ and its corresponding iterates variables $y_{t}$,
    $\mathbb{E}[\gamma y_{t}^{2}/2]\geq\mathbb{E}[\gamma(y_{t}^{*})^{2}/2]=\frac{\sigma^{2}}{\gamma}\eta_{t}^{*}.$
Proof.
We will denote $\sigma_{t}=\mathbb{E}[y_{t}^{2}]$ and assume WLOG we start decayping at step 00. Then we will have
$\sigma_{t}=(1-\eta_{t}\gamma)^{2}\sigma_{t-1}+\eta_{t}^{2}\sigma^{2}.$
If we choose all $\eta_{t}=\eta_{\mathrm{max}}$, we can directly verify that $\sigma_{t}=\sigma$.
If we choose $\eta_{t}=\frac{\sigma_{t-1}\gamma}{\sigma_{t-1}\gamma^{2}+\sigma^{2}}$ to minimize the right hand side, we will have that
$\sigma_{t}=\frac{\sigma_{t-1}\sigma^{2}}{\sigma_{t-1}\gamma^{2}+\sigma^{2}}.$
$\iff$ $\frac{1}{\sigma_{t}}=\frac{1}{\sigma_{t-1}}+\frac{\gamma^{2}}{\sigma^{2}}=\frac{1}{\sigma_{0}}+\frac{\gamma^{2}t}{\sigma^{2}}.$
This implies $\sigma_{t}=\frac{1}{\frac{1}{\sigma_{0}}+\frac{\gamma^{2}t}{\sigma^{2}}}$ and plugging into $\eta_{t}=\frac{\sigma_{t-1}\gamma}{\sigma_{t-1}\gamma^{2}+\sigma^{2}}$ we have that
$\eta_{t}^{*}=\frac{\gamma}{\gamma^{2}+\frac{\sigma^{2}}{\sigma_{t-1}}}=\frac{\gamma}{\gamma^{2}t+\frac{\sigma^{2}}{\sigma_{0}}}=\frac{\gamma}{\sigma^{2}}\sigma_{t}.$
The optimality of $\eta_{t}^{*}$ can be easily inferred from the proof. ∎
A.3 Landscape Analysis
We will parameterize ${P}_{F}(x)$ as $v_{d}(x)v_{d}(x)^{T}$ and ${P}_{S}(x)$ to denote $\mathcal{I}-{P}_{F}(x)$. Throughout this section, we will assume $v_{d}(x)$ is continuous and pointing towards the direction of the gradient for all the $x$ on the river. The following technical lemmas will be used repetitively in the proof.
Lemma A.2.
Under Assumptions 2 and 1, the directional derivative of ${P}_{S}(x)$ and ${P}_{F}(x)$ exist and satisfied that
$\nabla{P}_{F}(x)[v]=-\nabla{P}_{S}(x)[v]=\nabla v_{d}(x)[v]v_{d}(x)^{T}+v_{d}(x)\nabla v_{d}(x)[v]^{T}.$
Further
$\nabla{P}_{S}(x)[v]{P}_{S}a$ $=\langle\nabla v_{d}(x)[v],{P}_{S}a\rangle v_{d}(x),$
$\nabla{P}_{S}(x)[v]{P}_{F}a$ $=\langle v_{d},{P}_{F}a\rangle\nabla v_{d}(x)[v].$
$\|\nabla{P}_{S}(x)[v]\|_{2}$ $\leq\frac{\gamma\epsilon}{\Delta}\|v\|_{2}.$
Proof.
As $\gamma_{\mathrm{flat}}$ is an unique eigenvalue of $\nabla^{2}L(x+vt)$ and $\nabla^{2}L(x+vt)$ is analytical with respect to $t$, by Theorem 6.1 of Kato (1995), we know that $v_{d}(x+vt)$ is analytical with respect to $t$. Hence, the directional derivative exists.
The proof is by applying the chain rule and noticing that $\langle v_{d}(x),\nabla v_{d}(x)[v]\rangle=0$ because $v_{d}(x)$ is always a unit vector. ∎
We will now define the projection of iterate to the river as the progress measure of the optimization dynamics.
Definition A.3.
For $U$ in Assumption 2 and any $w\in U$, we define the following ODE as the projection flow:
$\phi(w,0)=w,d\phi(w,t)=-{P}_{S}(w)\nabla L\left(\phi(w,t)\right)dt.$ (9)
When $\lim_{t\to\infty}\phi(w,t)$ is well defined, we will define $\Phi(w)=\lim_{t\to\infty}\phi(w,t)$ as the projection of $w$ to the river.
The following lemma ensures that the projection function is well-defined and is close to $w$:
Lemma A.4.
Under Assumptions 1 and 2, for any $w$ satisfying that $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$, $\Phi(w)\in\mathcal{M}$ exists and $\|w-\Phi(w)\|_{2}\leq\frac{2\|{P}_{S}(w)\nabla L(w)\|_{2}}{\gamma+2\gamma_{\mathrm{flat}}}$. Moreover, movement along the projection flow decays exponentially, $\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|_{2}\leq\exp(-\gamma t/2)\|{P}_{S}(w)\nabla L(w)\|_{2}$.
Proof.
We will track $\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|_{2}^{2}$ along the projection flow before $\phi(w,t)$ leaves $U$,
$\frac{d\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|_{2}^{2}}{dt}$
$=$ $2\langle{P}_{S}(\phi(w,t))\nabla L(\phi(w,t)),\frac{d{P}_{S}(\phi(w,t))}{dt}\nabla L(\phi(w,t))+{P}_{S}(\phi(w,t))\frac{d\nabla L(\phi(w,t))}{dt}\rangle.$
By Lemmas A.2 and 2, the first term can be bounded as
$\langle{P}_{S}(\phi(w,t))\nabla L(\phi(w,t)),\frac{d{P}_{S}(\phi(w,t))}{dt}\nabla L(\phi(w,t))\rangle$
$=$ $-\langle{P}_{S}(\phi(w,t))\nabla L(\phi(w,t)),\nabla{P}_{S}(\phi(w,t))[{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))]\nabla L(\phi(w,t))\rangle$
$=$ $-\langle{P}_{S}(\phi(w,t))\nabla L(\phi(w,t)),\nabla v_{d}(x)[{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))]\rangle\langle v_{d},{P}_{F}(\phi(x,t))\nabla L(\phi(x,t))\rangle$
$\leq$ $\|\nabla L(\phi(w,t))\|\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|^{2}\frac{\kappa\gamma}{\Delta}\leq\kappa\gamma\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|^{2}.$
The second term is always negative
$\langle{P}_{S}(\phi(w,t))\nabla L(\phi(w,t)),{P}_{S}(\phi(w,t))\frac{d\nabla L(\phi(w,t))}{dt}\rangle$
$=$ $-\langle{P}_{S}(\phi(w,t))\nabla L(\phi(w,t)),{P}_{S}(\phi(w,t))\nabla^{2}L(\phi(x,t)){P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\rangle$
$\leq$ $-(\gamma+4\gamma_{\mathrm{flat}})\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|^{2}.$
Summing up the two terms,
$\frac{d\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|_{2}^{2}}{dt}\leq-(\gamma+2\gamma_{\mathrm{flat}})\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|^{2}.$
By Lemma A.34, we have $\|{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\|_{2}^{2}\leq\exp(-(\gamma+2\gamma_{\mathrm{flat}})t)\|{P}_{S}(w)\nabla L(w)\|_{2}^{2}$. Hence $\forall t>0$,
[…]
$\leq\|{P}_{S}(w)\nabla L(w)\|_{2}\int_{0}^{\infty}\exp(-(\gamma+2\gamma_{\mathrm{flat}})t/2)$
$\leq\frac{2\|{P}_{S}(w)\nabla L(w)\|_{2}}{(\gamma+2\gamma_{\mathrm{flat}})}.$
As $\mathcal{B}(w,\frac{\Delta}{2\gamma})\subset U$, the analysis hold along the trajectory and this shows that $\Phi(w)=\lim\limits_{t\to\infty}\phi(w,t)$ exists and that $\|\Phi(w)-w\|_{2}\leq\frac{2\|{P}_{S}(w)\nabla L(w)\|_{2}}{(\gamma+2\gamma_{\mathrm{flat}})}$.
Further $\Phi(w)$ satisfies that ${P}_{S}(\Phi(w))\nabla L(\Phi(w))=0$, and by Assumption 2, $\Phi(w)\in\mathcal{M}$. ∎
The following lemmas focus on the properties of $\partial\Phi$.
Lemma A.5.
Under Assumptions 1 and 2, for any $w$ satisfying that $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$, $\partial\Phi(w)$ is well-defined.
Proof.
Recall that $\Phi(w)=\lim\limits_{n\to\infty}\underbrace{(\phi\circ\phi\circ\ldots\phi)}_{n\text{ times}}(w)$, as $\phi$ is differentiable (Lemma A.2) and $\mathcal{M}$ is the fixed point of $\Phi$, by Theorem 5.1 of Falconer (1983), we have that $\partial\Phi(w)$ is well-defined. ∎
Lemma A.6.
Under Assumptions 1 and 2, for any $w$ satisfying that $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$, it holds that $\partial\Phi(w){P}_{S}(w)\nabla L(w)=0$. Further for any $w=x(t)\in\mathcal{M}$, it holds that $\partial\Phi(w){P}_{S}(w)=0,\partial\Phi(w)\frac{dx(t)}{dt}=\frac{dx(t)}{dt}$ and that for any $v$, $\partial\Phi(w)v$ aligns with $\frac{dx(t)}{dt}$.
Proof.
According to Lemmas A.4 and A.6, $\Phi,\partial\Phi$ is well-defined when $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$. Based on Definition A.3, we have that
$\forall t,\Phi(\phi(w,t))=\Phi(w).$
Hence,
$\frac{d\Phi(\phi(w,t))}{dt}\mid_{t=0}=0,$
Therefore,
$0=\partial\Phi(w)\frac{d\phi(w,t)}{dt}\mid_{t=0}=-\partial\Phi(w){P}_{S}(w)\nabla L(w).$
For any $w\in\mathcal{M}$ and any $v\in\mathbb{R}^{d}$, it holds that
00 0 $=\frac{d\partial\Phi(w+\alpha v){P}_{S}(w+\alpha v)\nabla L(w+\alpha v)}{d\alpha}\mid_{\alpha=0}$
$=\partial^{2}\Phi(w)[v]{P}_{S}(w)\nabla L(w)+\partial\Phi(w)\Big{[}\partial{P}_{S}(w)[v]\nabla L(w)+{P}_{S}(w)\nabla^{2}L(w)v\Big{]}$
$=\partial\Phi(w)\Big{[}\partial{P}_{S}(w)[v]\nabla L(w)+{P}_{S}(w)\nabla^{2}L(w)v\Big{]}.$ (10)
Define $J_{w}(v)$ as the projection from $v$ to $\partial{P}_{S}(w)[v]\nabla L(w)+{P}_{S}(w)\nabla^{2}L(w)v$.
Lemma A.7.
$J_{w}(v)$ is a linear projection and the range of $J_{w}$ is the range of ${P}_{S}$.
Proof.
Based on Lemma A.2, ${P}_{S}\partial{P}_{S}(w)[v]\nabla L(w)=\partial{P}_{S}(w)[v]\nabla L(w)$. Hence the range of $J_{w}$ is a subspace of the range of ${P}_{S}$. When $v={P}_{S}(w)u\neq 0$, based on Assumption 2,
$\|J_{w}(v)\|_{2}\geq\|{P}_{S}\nabla^{2}L(w){P}_{S}(w)u\|_{2}-\|\partial{P}_{S}(w)[{P}_{S}(w)u]\nabla L(w)\|_{2}\geq\gamma\|u\|_{2}-\gamma\kappa\|u\|_{2}>0.$
Hence the range of $J_{w}$ has a dimension no smaller than the dimension of the range of ${P}_{S}(w)$. This concludes that the range of $J_{w}$ is the range of ${P}_{S}(w)$. ∎
Hence by Equations 10 and A.7, it holds that for $w\in\mathcal{M}$, $\partial\Phi(w){P}_{S}(w)=0$. This shows that the range of $\partial\Phi(w)$ has dimension $1$.
Finally for any $w\in\mathcal{M}$, $\Phi(w)=w$. Hence,
$\frac{d\Phi(x(t))-dx(t)}{dt}=0.$
$\partial\Phi(x(t))\frac{dx(t)}{dt}=\frac{dx(t)}{dt}.$
Hence the range of $\partial\Phi(w)$ contains $\frac{dx(t)}{dt}$, this concludes the proof. ∎
Lemma A.8.
Under Assumptions 1 and 2, for any $w$ satisfying that $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$, it holds that
$\frac{\|{P}_{F}[\Phi(w)]\partial\Phi(w)\nabla L(w)-{P}_{F}(w)\nabla L(w)\|_{2}}{\|{P}_{F}(w)\nabla L(w)\|_{2}}$ $\leq 5\kappa,$
$\frac{\|{P}_{S}[\Phi(w)]\partial\Phi(w)\nabla L(w)\|_{2}}{\|{P}_{F}(w)\nabla L(w)\|_{2}}$ $\leq 5\kappa.$
Proof.
First, by Lemma A.6, it holds that $\partial\Phi(w)\nabla L(w)=\partial\Phi(w){P}_{F}(w)\nabla L(w)$. Define
$v$ $={P}_{F}(w)\nabla L(w)/\|{P}_{F}(w)\nabla L(w)\|_{2},$
$s(t)$ $={P}_{S}(\phi(w,t))\partial\phi(w,t)[v],$
$f(t)$ $={P}_{F}(\phi(w,t))\partial\phi(w,t)[v],$
it holds that $s(0)=0$ and $f(0)=v$ as $\phi(w,0)=w$.
We will bound the changes of $s(t)$ and $f(t)$. We will begin with calculating the time derivative of $\partial\phi(w,t)[v]$.
$\frac{d\partial\phi(w,t)[v]}{dt}=$ $\partial\left(\frac{d\phi(w,t)}{dt}\right)[v]$
$=$ $-\partial\left({P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\right)[v]$
$=$ $-\partial{P}_{S}(\phi(w,t))[\partial\phi(w,t)[v]]\nabla L(\phi(w,t))$
$-{P}_{S}(\phi(w,t))\nabla^{2}L(\phi(w,t))\partial\phi(w,t)[v]$
$=$ $-\partial{P}_{S}(\phi(w,t))[\partial\phi(w,t)[v]]\nabla L(\phi(w,t))$
$-{P}_{S}(\phi(w,t))\nabla^{2}L(\phi(w,t))s(t).$ (11)
We will now bound $\frac{d\|s(t)\|_{2}}{dt}$,
$\frac{d\|s(t)\|_{2}^{2}}{dt}=$ $2\langle s(t),\frac{ds(t)}{dt}\rangle$
$=$ $2\langle s(t),\nabla{P}_{S}(\phi(w,t))[\frac{d\phi(w,t)}{dt}]\partial\phi(w,t)[v]+{P}_{S}(\phi(w,t))\frac{d\partial\phi(w,t)[v]}{dt}\rangle$
$=$ $-2\langle s(t),\nabla{P}_{S}(\phi(w,t))[{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))]\partial\phi(w,t)[v]\rangle$
$-2\langle s(t),{P}_{S}(\phi(w,t))\frac{d\partial\phi(w,t)[v]}{dt}\rangle.$
By Lemmas A.4, 2 and A.2, the first term satisfies that,
$\langle s(t),\nabla{P}_{S}(\phi(w,t))[{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))]\partial\phi(w,t)[v]\rangle$ $\leq\gamma\kappa\|s(t)\|_{2}\|s(t)+f(t)\|_{2}.$
By Equations 11 and 2, the second term satisfies that,
$\langle s(t),{P}_{S}(\phi(w,t))\frac{d\partial\phi(w,t)[v]}{dt}\rangle$
$=$ $-\langle s(t),\partial{P}_{S}(\phi(w,t))[\partial\phi(w,t)[v]]\nabla L(\phi(w,t))\rangle$
$-\langle s(t),{P}_{S}(\phi(w,t))\nabla^{2}L(\phi(w,t))s(t)\rangle$
$\leq$ $\gamma\kappa\|s(t)\|_{2}\|s(t)+f(t)\|_{2}-\gamma\|s(t)\|_{2}^{2}.$
Hence $2\|s(t)\|_{2}\frac{d\|s(t)\|_{2}}{dt}=\frac{d\|s(t)\|_{2}^{2}}{dt}\leq-2\gamma\|s(t)\|_{2}^{2}+4\gamma\kappa\|s(t)\|_{2}\|s(t)+f(t)\|_{2}$ and we can conclude that
$\frac{d\|s(t)\|_{2}}{dt}\leq-\gamma\|s(t)\|_{2}/2+2\gamma\kappa\|f(t)\|_{2}.$ (12)
Similarly, we can provide a bound for $\|\frac{df(t)}{dt}\|_{2}$,
$\|\frac{df(t)}{dt}\|=$ $2\|\nabla{P}_{F}(\phi(w,t))[\frac{d\phi(w,t)}{dt}]\partial\phi(w,t)[v]+{P}_{F}(\phi(w,t))\frac{d\partial\phi(w,t)[v]}{dt}\|_{2}$
$=$ $-2\|\nabla{P}_{F}(\phi(w,t))[{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))]\partial\phi(w,t)[v]$
$+{P}_{F}(\phi(w,t))\frac{d\partial\phi(w,t)[v]}{dt}\|_{2}.$
By Lemmas A.4, 2 and A.2, the first term satisfies that,
$\|\nabla{P}_{F}(\phi(w,t))[{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))]\partial\phi(w,t)[v]\|$
$=$ $\|\nabla{P}_{S}(\phi(w,t))[{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))]\partial\phi(w,t)[v]\|$
$\leq$ $\gamma\kappa\exp(-\gamma t/2)\|f(t)\|_{2}\|s(t)+f(t)\|_{2}.$
By Lemmas A.4, 2 and A.2, the second term satisfies that,
$\|{P}_{F}(\phi(w,t))\frac{d\partial\phi(w,t)[v]}{dt}\|$
$=$ $\|{P}_{F}(\phi(w,t))\partial{P}_{F}(\phi(w,t))[\partial\phi(w,t)[v]]\nabla L(\phi(w,t))\rangle\|$
$=$ $\|{P}_{F}(\phi(w,t))\partial{P}_{F}(\phi(w,t))[\partial\phi(w,t)[v]]{P}_{S}(\phi(w,t))\nabla L(\phi(w,t))\rangle\|$
$\leq$ $\gamma\kappa\exp(-\gamma t/2)\|f(t)\|_{2}\|s(t)+f(t)\|_{2}.$
Hence we can conclude that
$\|\frac{df(t)}{dt}\|_{2}\leq 2\gamma\kappa\exp(-\gamma t/2)(\|f(t)\|_{2}+\|s(t)\|_{2}).$ (13)
By Equation 12, it holds that
$\frac{d(\exp(\gamma t/2)\|s(t)\|_{2})}{dt}=\gamma\exp(\gamma t/2)\|s(t)\|_{2}/2+\exp(\gamma t/2)\frac{d\|s(t)\|_{2}}{dt}\leq 2\gamma\kappa\exp(\gamma t/2)\|f(t)\|_{2}.$
Integrating the above equation from 00 to $t$, and we have
[…]
By Equations 13 and 14, we have that
[…]
This suggests that
[…]
Define $M(t)=\sup_{0\leq\tau\leq t}f(t)$, then it holds that
[…]
$\leq 1+\|M(T)\|_{2}\left(4\kappa+16\kappa^{2}\right).$
This implies that $\forall t,\|f(t)\|_{2}\leq\|M(t)\|_{2}\leq\frac{1}{1-4\kappa-16\kappa^{2}}\leq 1+5\kappa$. By Equation 14, this suggests that $\|s(t)\|_{2}\leq 4\kappa(1+5\kappa)\leq 5\kappa$. Finally, returning to Equation 13, we have that
$\|\frac{f(t)}{dt}\|_{2}$ $\leq 2\gamma\kappa\exp(-\gamma t/2)(\|f(t)\|_{2}+\|s(t)\|_{2})$
$\leq 2\gamma\kappa\exp(-\gamma t/2)(1+10\kappa).$
Hence $\|f(t)-f(0)\|_{2}\leq 2\int_{0}^{\infty}2\gamma\kappa\exp(-\gamma t/2)(1+10\kappa)\leq 2\kappa(1+10\kappa)\leq 5\kappa$.
We have that
$\frac{\|{P}_{F}[\Phi(w)]\partial\Phi(w)\nabla L(w)-{P}_{F}(w)\nabla L(w)\|_{2}}{\|{P}_{F}(w)\nabla L(w)\|_{2}}=\lim_{t\to\infty}\|f(t)-f(0)\|_{2}\in[0,5\kappa],$
and
$\frac{\|{P}_{S}[\Phi(w)]\partial\Phi(w)\nabla L(w)\|_{2}}{\|{P}_{F}(w)\nabla L(w)\|_{2}}=\lim_{t\to\infty}\|s(t)\|_{2}\in[0,5\kappa],$
The proof is then complete. ∎
The following lemma generalizes Lemma A.8 to general direction instead of $\nabla L(w)$.
Lemma A.9.
Under Assumptions 1 and 2, for any $w$ satisfying that $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$, it holds that
$\frac{\|{P}_{F}[\Phi(w)]\partial\Phi(w)u-{P}_{F}(w)u\|_{2}}{\|{P}_{F}(w)\nabla L(w)\|_{2}}$ $\leq 5\kappa,$
$\frac{\|{P}_{S}[\Phi(w)]\partial\Phi(w)u\|_{2}}{\|{P}_{F}(w)u\|_{2}}$ $\leq 5\kappa.$
Proof.
We only need to notice that ${P}_{F}(w)u$ aligns with $v_{d}(\nabla^{2}L(w))$. Hence, it always holds that
${P}_{F}(w)u$ $={P}_{F}(w)\nabla L(w)\frac{\langle{P}_{F}(w)\nabla L(w),{P}_{F}(w)u\rangle}{\|{P}_{F}(w)\nabla L(w)\|_{2}^{2}},$
$\partial\Phi(w)u$ $=\partial\Phi(w){P}_{F}(w)u=\partial\Phi(w)\nabla L(w)\frac{\langle{P}_{F}(w)\nabla L(w),{P}_{F}(w)u\rangle}{\|{P}_{F}(w)\nabla L(w)\|_{2}^{2}}.$
The proof is then complete. ∎
The following lemma states that the angle between the gradient and the tangent direction is small for any point on the river.
Lemma A.10.
For any $w\in\mathcal{M}$, it holds that
$\|{P}_{\mathcal{M}}(w)\nabla L(w)-\nabla L(w)\|_{2}\leq 4\kappa\|{P}_{\mathcal{M}}(w)\nabla L(w)\|_{2}.$
Proof.
Assume $w=x(T)$, we will denote ${P}_{\mathcal{M}}(w)\nabla L(w)$ by $v$.
It holds that
$\nabla\big{(}{P}_{S}(w)\nabla L(w)\big{)}[v]=0,$
which can be simplified to
${P}_{S}(w)\nabla^{2}L(w)v+\nabla{P}_{S}(w)[v]\nabla L(w)=0.$
The first term satisfies that $\|{P}_{S}(w)\nabla^{2}L(w)v\|_{2}\geq\gamma\|{P}_{S}(w)v\|$ and the second term satisfies that $\|\nabla{P}_{S}(w)[v]\nabla L(w)\|_{2}\leq\gamma\kappa\|v\|$. This then suggests $\|{P}_{S}(w)v\|_{2}\leq\kappa\|v\|_{2}$.
Therefore $\|{P}_{F}(w)v\|_{2}\geq(1-\kappa)\|v\|_{2}$. As
$v=\frac{dx(t)}{dt}\mid_{t=T}=-{P}_{\mathcal{M}}\left(w\right)\nabla L(w)$
We know that $\Big{|}v_{d}^{\top}{P}_{\mathcal{M}}(w)v_{d}\Big{|}\geq(1-\kappa)\|{P}_{\mathcal{M}}\left(w\right)v_{d}\|_{2}$, which suggests that $\Big{|}v_{d}^{\top}\frac{{P}_{\mathcal{M}}(w)v_{d}}{\|{P}_{\mathcal{M}}(w)v_{d}\|_{2}}\Big{|}\geq(1-\kappa)$. Hence we can conclude that $\|{P}_{\mathcal{M}}(w)v_{d}\|_{2}\geq(1-\kappa)$. Hence, we know that
$\|v+\nabla L(w)\|_{2}\leq\sqrt{1-(1-\kappa)^{2}}\|\nabla L(w)\|_{2}\leq 2\kappa\|\nabla L(w)\|_{2}\leq\frac{2\kappa}{1-\kappa}\|v\|_{2}\leq 4\kappa\|v\|_{2}.$
This concludes the proof. ∎
The next lemma states that ${P}_{F}(w)\nabla L(w)$ and $\nabla L(\Phi(w))$ is always close.
Lemma A.11.
Under Assumptions 1 and 2, for any $w$ satisfying that $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$, it holds that
$\|{P}_{F}(w)\nabla L(w)-\nabla L(\Phi(w))\|_{2}\leq(\gamma\kappa+\gamma_{\mathrm{flat}})\|w-\Phi(w)\|_{2}.$
Proof.
By Lemma A.4, the line segment from $\Phi(w)$ to $w$ lies in $U$. By Assumptions 2 and A.4,
$\|{P}_{F}(w)\nabla L(w)-{P}_{F}(\Phi(w))\nabla L(\Phi(w))\|_{2}$
$=$ $\|\int_{0}^{1}\nabla{P}_{F}\left(\Phi\left(w\right)+t\left(w-\Phi\left(w\right)\right)\right)\left[w-\Phi(w)\right]\nabla L(\Phi\left(w\right)+t\left(w-\Phi\left(w\right)\right))dt$
[…]
$\leq$ $(\gamma\kappa+\gamma_{\mathrm{flat}})\|w-\Phi(w)\|_{2}.$
This concludes the proof. ∎
The final theorem states that when $w$ is near the river, the movement of its projection has a similar value as the inherent speed at the river.
Lemma A.12.
Under Assumptions 1 and 2, when $\|w-\Phi(w)\|_{2}\leq\frac{10\kappa\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}}{\gamma+\gamma_{\mathrm{flat}}}$,
$\|{P}_{F}(w)\nabla L(w)+\frac{dx(\tau)}{d\tau}\mid_{\tau=T}\|_{2}$ $\leq 16\kappa\|\frac{dx(\tau)}{d\tau}\mid_{\tau=T}\|_{2}.$
$\|\partial\Phi(w)\nabla L(w)+\frac{dx(\tau)}{d\tau}\mid_{\tau=T}\|_{2}$ $\leq 30\kappa\|\frac{dx(\tau)}{d\tau}\mid_{\tau=T}\|_{2}.$
Proof.
By Lemma A.8
$\|\partial\Phi(w)\nabla L(w)-{P}_{F}(w)\nabla L(w)\|_{2}\leq 10\kappa\|{P}_{F}(w)\nabla L(w)\|_{2}.$
Combining Lemma A.11 and $\|w-\Phi(w)\|_{2}\leq\frac{10\kappa\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}}{\gamma+\gamma_{\mathrm{flat}}}$, we have that
$\|{P}_{F}(w)\nabla L(w)-\nabla L(\Phi(w))\|_{2}\leq(\gamma\kappa+\gamma_{\mathrm{flat}})\|w-\Phi(w)\|\leq 10\kappa\|{P}_{F}(w)\nabla L(w)\|_{2}.$
By Lemma A.10, let $v=\frac{dx(\tau)}{d\tau}\mid_{\tau=T}$,
$\|v+\nabla L(\Phi(w))\|_{2}\leq 4\kappa\|v\|_{2}.$
Combining the three inequalities, we have that
$\|{P}_{F}(w)\nabla L(w)-\nabla L(\Phi(w))\|_{2}$ $\leq\frac{10\kappa}{1-10\kappa}\|\nabla L(\Phi(w))\|_{2}\leq\frac{1+4\kappa}{1-10\kappa}10\kappa\|v\|_{2}\leq 12\kappa\|v\|_{2}.$
This suggests that
$\|{P}_{F}(w)\nabla L(w)-v\|_{2}\leq|{P}_{F}(w)\nabla L(w)-\nabla L(\Phi(w))\|_{2}+\|v+\nabla L(\Phi(w))\|_{2}\leq 16\kappa\|v\|_{2}.$
Hence
$\|v+\partial\Phi(w)\nabla L(w)\|_{2}$
$\leq$ $\|{P}_{F}(w)\nabla L(w)-v\|_{2}+\|{P}_{F}(w)\nabla L(w)-\nabla L(\Phi(w))\|_{2}$
$\leq$ $16\kappa\|v\|_{2}+10\kappa\|{P}_{F}(w)\nabla L(w)\|_{2}$
$\leq$ $30\kappa\|v\|_{2}.$
This concludes the proof. ∎
A.4 Proof of Theorem 3.2
We will consider the following gradient flow:
$dw(t)=-\nabla L(w(t))dt,w(0)\in V.$ (15)
We will first prove that along the gradient flow trajectory, it holds that $\|{P}_{S}(w)\nabla L(w)\|_{2}$ is bounded.
Lemma A.13.
Under Assumptions 1 and 2, along the gradient flow Equation 15, it holds that for $t\geq{2\log(2\Delta/(\kappa\Delta_{\min}))}/{\gamma}$, $\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}\leq 2\kappa\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}$.
Proof.
We will first compute how fast $\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}$ can change
$\frac{d\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}^{2}}{dt}=$ $-2\langle{P}_{S}(w(t))\nabla L(w(t)),\partial{P}_{S}(w(t))[\nabla L(w(t))]\nabla L(w(t))\rangle$
$-2\langle{P}_{S}(w(t))\nabla L(w(t)),{P}_{S}(w(t))\nabla^{2}L(w(t))\nabla L(w(t))\rangle$
By Lemmas A.2 and 2, the first term satisfies,
$-2\langle{P}_{S}(w(t))\nabla L(w(t)),\partial{P}_{S}(w(t))[\nabla L(w(t))]\nabla L(w(t))\rangle$
$=$ $-2\langle{P}_{S}(w(t))\nabla L(w(t)),\partial{P}_{S}(w(t))[\nabla L(w(t))]{P}_{F}L(w(t))\nabla L(w(t))\rangle$
$\leq$ $2\kappa\gamma\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}$
By Assumption 2, the second term satisfies,
$-2\langle{P}_{S}(w(t))\nabla L(w(t)),{P}_{S}(w(t))\nabla^{2}L(w(t))\nabla L(w(t))\rangle$
$\leq-2(\gamma+\gamma_{\mathrm{flat}})\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}^{2}.$
Hence,
$\frac{d\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}}{dt}\leq\kappa\gamma\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}-(\gamma+\gamma_{\mathrm{flat}})\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}.$ (16)
We then consider the corresponding ${P}_{F}(w(t))\nabla L(w(t))$.
$\frac{d\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}^{2}}{dt}=$ $-2\langle{P}_{F}(w(t))\nabla L(w(t)),\partial{P}_{F}(w(t))[\nabla L(w(t))]\nabla L(w(t))\rangle$
$-2\langle{P}_{F}(w(t))\nabla L(w(t)),{P}_{F}(w(t))\nabla^{2}L(w(t))\nabla L(w(t))\rangle$
By Lemmas A.2 and 2, the first term satisfies,
$-2\langle{P}_{F}(w(t))\nabla L(w(t)),\partial{P}_{F}(w(t))[\nabla L(w(t))]\nabla L(w(t))\rangle$
$=$ $-2\langle{P}_{F}(w(t))\nabla L(w(t)),\partial{P}_{F}(w(t))[\nabla L(w(t))]{P}_{S}L(w(t))\nabla L(w(t))\rangle$
$\leq$ $2\kappa\gamma\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}$
By Assumption 2, the second term satisfies,
$-2\langle{P}_{F}(w(t))\nabla L(w(t)),{P}_{F}(w(t))\nabla^{2}L(w(t))\nabla L(w(t))\rangle$
$\leq 2\gamma_{\mathrm{flat}}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}^{2}.$
Hence, we have that
$\Big{|}\frac{d\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}}{dt}\Big{|}\leq\kappa\gamma\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}+\gamma_{\mathrm{flat}}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}.$ (17)
Choose $\alpha_{\kappa}=\frac{1-\sqrt{1-4\kappa^{2}}}{2\kappa}<1.5\kappa$ as the solution to the quadratic equation $\kappa\alpha^{2}-\alpha+\kappa=0$.
Then combining Equations 16 and 17, it holds that
$\frac{d\left(\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}-\alpha_{\kappa}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}\right)}{dt}$
$\leq$ $\gamma(-1+\kappa\alpha_{\kappa})\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}+\kappa\gamma\|{P}_{F}\|_{2}$
$-\gamma_{\mathrm{flat}}\left(|{P}_{S}(w(t))\nabla L(w(t))\|_{2}-\alpha_{\kappa}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}\right).$
Notice that
$\frac{(-1+\kappa\alpha_{\kappa})}{\kappa}=\frac{-1}{\alpha_{\kappa}}$
Hence,
$\frac{d\left(\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}-\alpha_{\kappa}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}\right)}{dt}$
$\leq$ $-(\gamma(1-\kappa\alpha_{\kappa})+\gamma_{\mathrm{flat}})\left(\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}-\alpha_{\kappa}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}\right).$
By Lemma A.34, this suggests that
$\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}-\alpha_{\kappa}\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}$
$\leq$ $\exp(-(\gamma(1-\kappa\alpha_{\kappa})+\gamma_{\mathrm{flat}})t)\left(\|{P}_{S}(w(0))\nabla L(w(0))\|_{2}\right)$
$\leq$ $\exp(-\gamma t/2)\|{P}_{S}(w(0))\nabla L(w(0))\|_{2}.$
Hence,
$\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}\leq 1.5\kappa\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}+\exp(-\gamma t/2)\left(\|{P}_{S}(w(0))\nabla L(w(0))\|_{2}\right).$
Lemma A.14.
Under Assumptions 1 and 2, along the gradient flow Equation 15, it holds that $w(t)\in U$ and $\|w(t)-\Phi(w(t))\|_{2}\leq\frac{4\kappa\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}+2\exp(-\gamma t/2)\Delta}{\gamma+\gamma_{\mathrm{flat}}}$.
Proof.
This is a direct combination of Lemmas A.13 and A.4. ∎
Lemma A.15.
Under Assumptions 1 and 2, along the gradient flow Equation 15, if $T(t)$ satisfies $x(T(t))=\Phi(w(t))$, then
$\frac{dT(t)}{dt}\in[1-30\kappa,1+30\kappa].$
Proof.
As $T(t)$ satisfies $x(T(t))=\Phi(w(t))$, taking derivative on both sides yield,
$\frac{dT(t)}{dt}\frac{dx(\tau)}{d\tau}\mid_{\tau=T(t)}=-\partial\Phi(w(t))\nabla L(w(t)).$
By Lemmas A.14 and A.11, it holds that
$\|\partial\Phi(w(t))\nabla L(w(t))-\frac{dx(\tau)}{d\tau}\mid_{\tau=T(t)}\|_{2}\leq 30\kappa\frac{dx(\tau)}{d\tau}\mid_{\tau=T(t)}.$
We then have that
$|\frac{dT(t)}{dt}-1|\leq 30\kappa,$
which concludes the proof. ∎
Proof of Theorem 3.2.
The proof is a direct combination of Lemmas A.14 and A.15. ∎
A.5 Proof of Theorem 3.3
We will consider the following gradient descent:
$w_{t+1}-w_{t}=-\eta\nabla L(w_{t}),w_{0}\in\mathcal{M}.$ (18)
We will track the changes of ${P}_{F}(w(t))\nabla L(w(t))$ and ${P}_{S}(w(t))\nabla L(w(t))$, for simplicity, we will denote them us $fg(t)$ and $sg(t)$. Further, we will use the following denotation
$w_{t,\tau}=(1-\tau)w_{t}+\tau w_{t+1}$
We will first prove some lemmas bounding the difference between gradient and projections at different points.
Lemma A.16.
Under Assumptions 1 and 2, when $w_{t}\in V$, $\forall\tau\in(0,1)$, $w_{t,\tau}\in U$.
Proof.
It holds that,
$\|w_{t}-w_{t,\tau}\|_{2}\leq\eta\Delta\leq\frac{\Delta}{2\gamma}.$
Lemma A.17.
Under Assumptions 1 and 2, when $w_{t}\in V$, $\forall\tau,\tau^{\prime}\in[0,1]$, it holds that
$\|{P}_{S}(w_{t,\tau})-{P}_{S}(w_{t,\tau^{\prime}})\|_{2}\leq\eta\gamma\kappa.$
Proof.
According to Lemma A.16, it holds that $w_{t,\tau},w_{t,\tau^{\prime}}\in U$. Assume without loss of generality $\tau>\tau^{\prime}$,
[…]
$\leq$ $\eta\gamma\kappa.$
Lemma A.18.
Under Assumptions 1 and 2, when $w_{t}\in V$, $\forall\tau\in(0,1)$, it holds that
$\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}\leq\eta\gamma_{\mathrm{max}}\|sg(t)\|_{2}+2\eta^{2}\gamma_{\mathrm{max}}\gamma\kappa\|\nabla L(w_{t})\|_{2}.$
Proof.
According to Lemma A.16, it holds that $w_{t,\tau},w_{t,\tau^{\prime}}\in U$. Define $g(\tau^{\prime})=\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau^{\prime}})-{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}$, then by Lagrange’s Mean Value Theorem, there exists $\tau^{\prime}$, such that
$\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}=g(\tau)-g(0)$
$=$ $\tau g^{\prime}(\tau^{\prime})=\tau\frac{d\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau^{\prime}})-{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}}{d\tau^{\prime}}$
$\leq$ $\|\frac{d{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau^{\prime}})-{P}_{S}(w_{t,\tau})\nabla L(w_{t})}{d\tau^{\prime}}\|_{2}$
$=$ $\eta\|{P}_{S}(w_{t,\tau})\nabla^{2}L(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}$
$\leq$ $\eta\|{P}_{S}(w_{t,\tau})\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{S}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}+\eta\|{P}_{S}(w_{t,\tau})\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{F}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}$
$\leq$ $\eta\gamma_{\mathrm{max}}\|{P}_{S}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}+\eta\|({P}_{S}(w_{t,\tau})-{P}_{S}(w_{t,\tau^{\prime}}))\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{F}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}.$
By Lemma A.17, it holds that
$\gamma_{\mathrm{max}}\|{P}_{S}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}\leq\gamma_{\mathrm{max}}\|sg(t)\|_{2}+\eta\gamma_{\mathrm{max}}\gamma\kappa\|\nabla L(w_{t})\|_{2}.$
$\|({P}_{S}(w_{t,\tau})-{P}_{S}(w_{t,\tau^{\prime}})\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{F}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}\leq\eta\gamma\gamma_{\mathrm{flat}}\kappa\|\nabla L(w_{t})\|_{2}.$
Summing up and the proof is complete. ∎
Lemma A.19.
$\forall\tau\in(0,1)$, it holds that
$\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau})-sg(t)\|_{2}\leq\|sg(t)\|_{2}+3\eta\gamma\kappa\|\nabla L(w_{t})\|_{2}.$
Proof.
This is a direct combination of Lemmas A.18 and A.17, with
$\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{S}(w_{t})\nabla L(w_{t})\|_{2}$
$\leq$ $\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}+\|({P}_{S}(w_{t,\tau})-{P}_{S}(w_{t}))\nabla L(w_{t})\|_{2}.$
The proof is then complete. ∎
Lemma A.20.
$\forall\tau\in(0,1)$, it holds that
$\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{F}(w_{t,\tau})\nabla L(w_{t})\|_{2}\leq\eta\gamma_{\mathrm{flat}}\|fg(t)\|_{2}+2\eta^{2}\gamma_{\mathrm{max}}\gamma\kappa\|\nabla L(w_{t})\|_{2}.$
Proof.
Define $g(\tau^{\prime})=\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau^{\prime}})-{P}_{F}(w_{t,\tau})\nabla L(w_{t})\|_{2}$, then by Lagrange’s Mean Value Theorem, there exists $\tau^{\prime}$, such that
$\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{F}(w_{t,\tau})\nabla L(w_{t})\|_{2}=g(\tau)-g(0)$
$=$ $\tau g^{\prime}(\tau^{\prime})=\tau\frac{d\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau^{\prime}})-{P}_{F}(w_{t,\tau})\nabla L(w_{t})\|_{2}}{d\tau^{\prime}}$
$\leq$ $\|\frac{d{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau^{\prime}})-{P}_{F}(w_{t,\tau})\nabla L(w_{t})}{d\tau^{\prime}}\|_{2}$
$=$ $\eta\|{P}_{F}(w_{t,\tau})\nabla^{2}L(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}$
$\leq$ $\eta\|{P}_{F}(w_{t,\tau})\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{F}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}+\eta\|{P}_{F}(w_{t,\tau})\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{S}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}$
$\leq$ $\eta\gamma_{\mathrm{flat}}\|{P}_{F}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}+\eta\|({P}_{F}(w_{t,\tau})-{P}_{F}(w_{t,\tau^{\prime}}))\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{S}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}.$
By Lemma A.17, it holds that
$\gamma_{\mathrm{flat}}\|{P}_{F}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}\leq\gamma_{\mathrm{flat}}\|fg(t)\|_{2}+\eta\gamma_{\mathrm{flat}}\gamma\kappa\|\nabla L(w_{t})\|_{2}.$
$\|({P}_{F}(w_{t,\tau})-{P}_{F}(w_{t,\tau^{\prime}})\nabla^{2}L(w_{t,\tau^{\prime}}){P}_{F}(w_{t,\tau^{\prime}})\nabla L(w_{t})\|_{2}\leq\eta\gamma\gamma_{\mathrm{max}}\kappa\|\nabla L(w_{t})\|_{2}.$
Summing up and the proof is complete. ∎
Lemma A.21.
$\forall\tau\in(0,1)$, it holds that
$\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})-fg(t)\|_{2}\leq\|fg(t)\|_{2}+3\eta\gamma\kappa\|\nabla L(w_{t})\|_{2}.$
Proof.
This is a direct combination of Lemmas A.20 and A.17, with
$\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{F}(w_{t})\nabla L(w_{t})\|_{2}$
$\leq$ $\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})-{P}_{F}(w_{t,\tau})\nabla L(w_{t})\|_{2}+\|({P}_{F}(w_{t,\tau})-{P}_{F}(w_{t}))\nabla L(w_{t})\|_{2}.$
The proof is then complete. ∎
We will prove a discrete version of Lemma A.13.
Lemma A.22.
Under Assumptions 1 and 2, when $\eta<1/\gamma_{\mathrm{max}}$, along the gradient flow Equation 18, it holds that $\|{P}_{S}(w(t))\nabla L(w(t))\|_{2}\leq 10\kappa\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}$ as long as $w(\tau)\in U,\forall\tau\leq t$.
Proof.
We will first consider $sg(t)$, By Lagrange’s Mean Value Theorem, there exists $\tau$, such that,
$\|sg(t+1)\|_{2}^{2}-\|sg(t)\|_{2}^{2}$
$=$ $\|{P}_{S}(w_{t,1})\nabla L(w_{t,1})\|_{2}^{2}-\|{P}_{S}(w_{t,0})\nabla L(w_{t,0})\|_{2}^{2}=\frac{d\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau})\|_{2}^{2}}{d\tau}$
$=$ $-\eta\langle{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau}),\partial{P}_{S}(w_{t,\tau})[\nabla L(w_{t})]\nabla L(w_{t,\tau})+{P}_{S}(w_{t,\tau})\nabla^{2}L(w_{t,\tau})\nabla L(w)\rangle$
The first term satisfies that
$-\eta\langle{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau}),\partial{P}_{S}(w_{t,\tau})[\nabla L(w_{t})]\nabla L(w_{t,\tau})\rangle$
$\leq$ $\eta\gamma\kappa\|\nabla L(w_{t})\|_{2}\|{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau})\|_{2}$
$\leq$ $\eta\gamma\kappa\|\nabla L(w_{t})\|_{2}(2\|sg(t)\|_{2}+3\eta\gamma\kappa\|\nabla L(w_{t})\|_{2}).$
The second term satisfies that
$-\eta\langle{P}_{S}(w_{t,\tau})\nabla L(w_{t,\tau}),{P}_{S}(w_{t,\tau})\nabla^{2}L(w_{t,\tau})\nabla L(w_{t})\rangle$
$=-\eta\langle{P}_{S}(w_{t,\tau})\nabla L(w_{t}),{P}_{S}(w_{t,\tau})\nabla^{2}L(w_{t,\tau})\nabla L(w_{t})\rangle+\langle{P}_{S}(w_{t,\tau})(\nabla L(w_{t})-\nabla L(w_{t,\tau})),\nabla^{2}L(w_{t,\tau})\nabla L(w_{t})\rangle$
$\leq-\eta(\gamma+4\gamma_{\mathrm{flat}})\|{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}^{2}+\eta\gamma_{\mathrm{max}}\|{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}\|{P}_{S}(w_{t,\tau})(\nabla L(w_{t})-\nabla L(w_{t,\tau}))\|_{2}$
As we have that $\|a-b\|^{2}\geq\frac{\|a\|^{2}}{2}-4\|b\|^{2}$, by Lemma A.17, it holds that
$-\|{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}^{2}$
$=$ $-\|{P}_{S}(w_{t})\nabla L(w_{t})+\left({P}_{S}(w_{t,\tau})-{P}_{S}(w_{t})\right)\nabla L(w_{t})\|_{2}^{2}$
$\leq$ $-\eta\gamma\frac{\|{P}_{S}(w_{t})\nabla L(w_{t})\|_{2}^{2}}{2}+4\eta\gamma\|\left({P}_{S}(w_{t,\tau})-{P}_{S}(w_{t})\right)\nabla L(w_{t})\|_{2}^{2}$
$\leq$ $-\frac{\|{P}_{S}(w_{t})\nabla L(w_{t})\|_{2}^{2}}{2}+4\eta\gamma(\eta\gamma\kappa\|\nabla L(w_{t})\|)^{2}.$
$=$ $-\frac{\|{P}_{S}(w_{t})\nabla L(w_{t})\|_{2}^{2}}{2}+4(\eta\gamma)^{2}(\kappa\|\nabla L(w_{t})\|)^{2}.$
$\leq$ $-\|sg(t)\|_{2}^{2}+2\eta\gamma\kappa^{2}\|\nabla L(w_{t})\|^{2}$
Hence
$-\eta(\gamma+4\gamma_{\mathrm{flat}})\|{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}^{2}$
$\leq$ $-\eta(\gamma+4\gamma_{\mathrm{flat}})\|sg(t)\|_{2}^{2}+2\eta^{2}(\gamma+4\gamma_{\mathrm{flat}})\gamma\kappa^{2}\|\nabla L(w_{t})\|^{2}$
$\leq$ $-\eta(\gamma+4\gamma_{\mathrm{flat}})\|sg(t)\|_{2}^{2}+2\eta\gamma\kappa^{2}\|\nabla L(w_{t})\|^{2}$
By Lemmas A.18 and A.17 and $\eta\gamma_{\mathrm{max}}^{2}\leq\gamma/2$, it holds that
$\eta\gamma_{\mathrm{max}}\|{P}_{S}(w_{t,\tau})\nabla L(w_{t})\|_{2}\|{P}_{S}(w_{t,\tau})(\nabla L(w_{t})-\nabla L(w_{t,\tau}))\|_{2}$
$\leq$ $\eta^{2}\gamma_{\mathrm{max}}^{2}(\|sg(t)\|_{2}+\eta\gamma\kappa\|\nabla L(w_{t})\|_{2})(\|sg(t)\|_{2}+2\eta\gamma\kappa\|\nabla L(w_{t})\|_{2})$
$\leq$ $\eta\gamma(\|sg(t)\|_{2}+\kappa\|\nabla L(w_{t})\|_{2}/2)(\|sg(t)\|_{2}+\kappa\|\nabla L(w_{t})\|_{2})/2.$
Hence, we can conclude that
$\|sg(t+1)\|_{2}^{2}-\|sg(t)\|_{2}^{2}$
$-\eta((\gamma+4\gamma_{\mathrm{flat}})\|sg(t)\|_{2}^{2}+4\eta\gamma\kappa^{2}\|\nabla L(w_{t})\|^{2}$
$+\eta\gamma(\|sg(t)\|_{2}+\kappa\|\nabla L(w_{t})\|_{2}/2)(\|sg(t)\|_{2}+\kappa\|\nabla L(w_{t})\|_{2})/2$
Let $b=\kappa\|\nabla L(w_{t})\|_{2}$ and $a=sg(t)$, as $b(2a+3b/2)-a^{2}+4b^{2}+\frac{1}{2}(a+\frac{b}{2})(a+b)\leq-\frac{a^{2}}{4}+10b^{2}$, it holds that
$\|sg(t+1)\|_{2}^{2}-\|sg(t)\|_{2}^{2}\leq$ $-\eta\gamma\frac{\|sg(t)\|_{2}^{2}}{4}+10\eta\gamma\kappa^{2}\|\nabla L(w_{t})\|^{2}.-4\eta\gamma_{\mathrm{flat}}\|sg(t)\|_{2}^{2}$ (19)
Similarly, we can control the $fg(t)$ changes. By Lagrange’s Mean Value Theorem, there exists $\tau^{\prime}$, such that,
$\|fg(t+1)\|_{2}^{2}-\|fg(t)\|_{2}^{2}$
$=$ $\|{P}_{F}(w_{t,1})\nabla L(w_{t,1})\|_{2}^{2}-\|{P}_{F}(w_{t,0})\nabla L(w_{t,0})\|_{2}^{2}=\frac{d\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})\|_{2}^{2}}{d\tau}$
$=$ $-\eta\langle{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau}),\partial{P}_{F}(w_{t,\tau})[\nabla L(w_{t})]\nabla L(w_{t,\tau})+{P}_{F}(w_{t,\tau})\nabla^{2}L(w_{t,\tau})\nabla L(w)\rangle$
The first term satisfies that
$\eta\langle{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau}),\partial{P}_{F}(w_{t,\tau})[\nabla L(w_{t})]\nabla L(w_{t,\tau})\rangle$
$\leq$ $\gamma\eta\kappa\|\nabla L(w_{t})\|_{2}\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})\|_{2}.$
$\leq$ $\gamma\eta\kappa\|\nabla L(w_{t})\|_{2}(\|fg(t)\|_{2}+\eta\gamma_{\mathrm{flat}}\|fg(t)\|_{2}+2\eta^{2}\gamma_{\mathrm{max}}\gamma\kappa\|\nabla L(w_{t})\|_{2}).$
$\leq$ $4\gamma\eta\kappa\|\nabla L(w_{t})\|_{2}^{2}.$
Similarly, the second term satisfies that
$\eta\langle{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau}),{P}_{F}(w_{t,\tau})\nabla^{2}L(w_{t,\tau})\nabla L(w)\rangle$
$\leq$ $\eta\gamma_{\mathrm{flat}}\|{P}_{F}(w_{t,\tau})\nabla L(w_{t})\|_{2}\|{P}_{F}(w_{t,\tau})\nabla L(w_{t,\tau})\|_{2}$
$\leq$ $\eta\gamma_{\mathrm{flat}}(\|fg(t)\|_{2}+\eta\gamma\kappa\|\nabla L(w_{t})\|_{2})(\|fg(t)\|_{2}+\eta\gamma_{\mathrm{flat}}\|fg(t)\|_{2}+2\eta^{2}\gamma_{\mathrm{max}}\gamma\kappa\|\nabla L(w_{t})\|_{2})$
$\leq$ $2\eta\gamma_{\mathrm{flat}}(\|fg(t)\|_{2}+\eta\gamma\kappa\|\nabla L(w_{t})\|_{2})^{2}$
$\leq$ $4\eta\gamma_{\mathrm{flat}}\|fg(t)\|_{2}^{2}+4\eta^{2}\gamma^{2}\kappa^{2}\|\nabla L(w_{t})\|_{2}^{2}$
Summarizing and we have
$\|fg(t+1)\|_{2}^{2}-\|fg(t)\|_{2}^{2}\geq$ $-5\gamma\eta\kappa\|\nabla L(w_{t})\|_{2}^{2}-4\eta\gamma_{\mathrm{flat}}\|fg(t)\|_{2}^{2}$ (20)
Let $a_{\kappa}$ be the smaller positive solution of
$0\displaystyle 5\kappa a^{2}+(10\kappa^{2}+5\kappa-\frac{1}{4})a+10\kappa^{2}=$ 0 .
Then $a_{\kappa}=\frac{(-10\kappa^{2}-5\kappa+\frac{1}{4})-\sqrt{(-10\kappa^{2}-5\kappa+\frac{1}{4})^{2}-200\kappa^{3}}}{10\kappa}<100\kappa^{2}$.
Then combining Equations 20 and 19
$\|sg(t+1)\|_{2}^{2}-a_{\kappa}\|fg(t+1)\|_{2}^{2}$
$\leq$ $(1-4\eta\gamma_{\mathrm{flat}})(\|sg(t)\|_{2}^{2}-a_{\kappa}\|fg(t+1)\|_{2}^{2})-\eta\gamma(\frac{1}{4}+10\kappa^{2}-5\kappa a_{\kappa})\|sg(t)_{2}\|^{2}+\eta\gamma(10\kappa^{2}+5\kappa a_{\kappa})\|fg(t)_{2}\|^{2}$
$=$ $(1-4\eta\gamma_{\mathrm{flat}}-\eta\gamma(\frac{1}{4}+10\kappa^{2}-5\kappa a_{\kappa}))(\|sg(t)\|_{2}^{2}-a_{\kappa}\|fg(t)\|_{2}^{2}).$
As $\|sg(0)\|_{2}^{2}-a_{\kappa}\|fg(0)\|_{2}^{2}<0$, we have that $\|sg(t)\|_{2}^{2}<a_{\kappa}\|fg(t+1)\|_{2}^{2}<100\kappa^{2}\|fg(t)\|_{2}^{2}$ for all the $t$. ∎
Then we can show that gradient descent will also track the river closely.
Lemma A.23.
Under Assumptions 1 and 2, along the gradient flow Equation 15, it holds that $w(t)\in U$ and $\|w(t)-\Phi(w(t))\|_{2}\leq\frac{10\kappa\|{P}_{F}(w(t))\nabla L(w(t))\|_{2}}{\gamma+\gamma_{\mathrm{flat}}}$.
Proof.
This is a direct combination of Lemmas A.22 and A.4. ∎
Finally, we will show that the movement of the projection of the gradient flow moves approximately at the same rate as the river, a discrete version of Lemma A.14.
Lemma A.24.
Under Assumptions 1 and 2, along the gradient flow Equation 15, along the gradient descent Equation 18, if $T(t)$ satisfies $x(T(t))=\Phi(w_{[t],t-[t]})$ where $[t]$ is the integer part of $t$, then for any $t$ that is not integer
$\frac{dT(t)}{dt}\in[\eta-(30\kappa+4\eta\gamma_{\mathrm{flat}})\eta,\eta+(30\kappa+4\eta\gamma_{\mathrm{flat}})\eta].$
Proof.
Let $[t]=k,t-[t]=\tau$, as $T(t)$ satisfies $x(T(t))=\Phi(w(k,t-[t]))$, let $v=frac{dx(\tau)}{d\tau}\mid_{\tau=T(t)}$, taking derivative on both sides yield,
$\frac{dT(t)}{dt}v=-\eta\partial\Phi(w_{k,\tau})\nabla L(w_{k}).$
As the proof of Lemma A.20, there exists $\tau^{\prime}$
$\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})-{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})\|_{2}$
$\leq$ $\eta\|{P}_{F}(w_{k,\tau})\nabla^{2}L(w_{k,\tau^{\prime}}){P}_{F}(w_{k,\tau^{\prime}})\nabla L(w_{k})\|_{2}+\eta\|{P}_{F}(w_{k,\tau})\nabla^{2}L(w_{k,\tau^{\prime}}){P}_{S}(w_{k,\tau^{\prime}})\nabla L(w_{k})\|_{2}$
$\leq$ $\eta\gamma_{\mathrm{flat}}\|{P}_{F}(w_{k,\tau^{\prime}})\nabla L(w_{k})\|_{2}+\eta^{2}\gamma\kappa\gamma_{\mathrm{max}}\|\nabla L(w_{k})\|_{2}$
By Lemma A.17, it holds that
$\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})-{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})\|_{2}$
$\leq$ $\eta\gamma_{\mathrm{flat}}\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}+2\eta^{2}\gamma\gamma_{\mathrm{max}}\kappa\|\nabla L(w_{k})\|_{2}$
$\leq$ $\eta\gamma_{\mathrm{flat}}\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}+\kappa\|\nabla L(w_{k})\|_{2}$ (21)
By Lemma A.22,
$\|\nabla L(w_{k})\|_{2}$ $\leq\frac{1}{1-10\kappa}\|{P}_{F}(w_{k})\nabla L(w_{k})\|_{2}$
$\leq\frac{1}{1-10\kappa}\left(\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}+\eta\gamma\kappa\|\nabla L(w_{k})\|_{2}\right)$
This shows that
$\|\nabla L(w_{k})\|_{2}$ $\leq\frac{1}{1-10\kappa-\eta\gamma\kappa}\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}\leq(1+12\kappa)\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}$
Combining with Equation 21, we have that
$\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})-{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})\|_{2}$
$\leq$ $\eta\gamma_{\mathrm{flat}}\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}+\kappa(1+12\kappa)\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}$
$\leq$ $(\eta\gamma_{\mathrm{flat}}+2\kappa)\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}$
This shows that
$\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})-{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})\|_{2}\leq$ $\frac{(\eta\gamma_{\mathrm{flat}}+2\kappa)}{1-(\eta\gamma_{\mathrm{flat}}+2\kappa)}\|{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})\|_{2}$
$\leq$ $(2\eta\gamma_{\mathrm{flat}}+3\kappa)\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}$ (22)
By Lemma A.12
$\|{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})+v\|_{2}\leq 16\kappa\|v\|_{2}$ (23)
Combining Equations 23 and 22,
$\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})+v\|_{2}$
$\leq$ $\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})--{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})\|_{2}+\|{P}_{F}(w_{k,\tau})\nabla L(w_{k,\tau})+v\|_{2}$
$\leq$ $(2\eta\gamma_{\mathrm{flat}}+3\kappa)\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}+16\kappa\|v\|_{2}$
$\leq$ $((2\eta\gamma_{\mathrm{flat}}+3\kappa)(1+16\kappa)+16\kappa)\|v\|_{2}$
$\leq$ $(19\kappa+3\eta\gamma_{\mathrm{flat}})\|v\|_{2}.$ (24)
By Lemma A.9
$\|\partial\Phi(w_{k,\tau})\nabla L(w_{k})-{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}\leq 10\kappa\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}.$ (25)
Combining Equations 24 and 25, it holds that
$\|\partial\Phi(w_{k,\tau})\nabla L(w_{k})+v\|_{2}$
$\leq$ $\|\partial\Phi(w_{k,\tau})\nabla L(w_{k})-{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}+\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})+v\|_{2}$
$\leq$ $10\kappa\|{P}_{F}(w_{k,\tau})\nabla L(w_{k})\|_{2}+(19\kappa+3\eta\gamma_{\mathrm{flat}})\|v\|_{2}$
$\leq$ $(30\kappa+4\eta\gamma_{\mathrm{flat}})\|v\|_{2}.$
Hence
$\frac{dT(t)}{dt}\in[\eta-(30\kappa+4\eta\gamma_{\mathrm{flat}})\eta,\eta+(30\kappa+4\eta\gamma_{\mathrm{flat}})\eta].$
This concludes the proof. ∎
Proof of Theorem 3.3.
The proof is a direct combination of Lemmas A.23 and A.24. ∎
A.6 Proof of Theorem 3.4
We will first show that under Assumption 3, the loss is separable within $U$.
Lemma A.25.
Under Assumptions 1, 3 and 2, the river is a straight line parallel to $v_{d}$.
Proof.
In this case, the $\kappa$ in Assumption 2 is 00 and this is a direct corollary of of Lemma A.12. ∎
Lemma A.26.
Under Assumptions 1, 3 and 2, there exists functions $g$ and $h$, such that for any $w\in U$ satisfying that $\mathcal{B}(w,\frac{2\Delta}{\gamma})\subset U$, it holds that
$L(w)=g(\Phi(w))+h(w-\Phi(w)).$
Furthermore, $h$ is a $\gamma$-strongly convex function when constrained on the range of ${P}_{S}$.
Proof.
We will choose $g$ as the constraint of $L$ on $\mathcal{M}$. Now $w-\Phi(w)$ will always fall in the range of ${P}_{S}$. Consider any $y$ in the range of ${P}_{S}$ and as $\nabla v_{d}(w)[v]=0$, we have that
$y^{T}\nabla^{2}L(w)v_{d}=0.$
This then suggest that
$\nabla[\langle\nabla L(w),y\rangle][v_{d}]=0.$
We then have for any $a\in\mathcal{M}$, by Lemma A.25,
[…]
Now, as $h(w-\Phi(w))=L(w)-L(\Phi(w))$, $\nabla^{2}h(y)$ when constrained on the range of ${P}_{S}$ has an eigenvalue greater than $\gamma$. ∎
We will first consider the mixing dynamics of the current SGD iterates on a strongly convex loss $h$ with a minimizer at 00.
$y(t+1)=y(t)-\eta\nabla h(y(t))-\eta{g}_{t},y(0)=0,\mathbf{{g}_{t}}\sim\mathcal{N}(0,\sigma^{2}\mathcal{I})$ (26)
We will define a coupling process $\tilde{y}(t)$ as
$\tilde{y}(t+1)=\tilde{y}(t)-\eta H\tilde{y}(t)-\eta{g}_{t},w(0)=0,{g}_{t}\sim\mathcal{N}(0,\sigma^{2}\mathcal{I}),\tilde{y}(0)=0.$ (27)
Here $H=\nabla^{2}h(0)$ is positive definite.
Assumption 6 (Regularity of $h$).
We will assume the following for the function $h$, constant $\delta\in(0,1]$, learning rate $\eta$.
  1. 1.
    The smallest eigenvalue of $\nabla^{2}h(y)$ within $\mathcal{B}(0,r)$ is at least $\gamma>0$ and the largest eigenvalue for $H$ is at most $\gamma_{\mathrm{max}}$.
  2. 2.
    $\forall y,h(y)\in[0,M]$.
  3. 3.
    $\forall y\in\mathcal{B}(0,r),\|\nabla h(y)\|_{2}\leq\Delta,\|\nabla^{3}h(y)\|_{2}\leq\rho$.
  4. 4.
    $T>1/\gamma$.
  5. 5.
    $\eta<1/(2\gamma_{\mathrm{max}})$.
  6. 6.
    $\eta\rho^{2}\sigma^{2}\leq{\gamma^{3}}/{(1600d\log(8\gamma T/\delta))}$.
  7. 7.
    $10\frac{\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(8\gamma T/\delta)}+400\eta\rho\sigma^{2}d\log(8\gamma T/\delta)/\gamma^{2}\leq r$.
We will first show that $\tilde{y}(t)$ will be bounded with a high probability for $T/\eta$ steps.
Lemma A.27.
For any $\delta\in(0,1]$, with probability $1-\delta$, for $\tilde{y}_{t}$ defined in Equation 27, under Assumption 6, it holds that for any $t\leq T/\eta$,
$\|\tilde{y}_{t}\|_{2}\leq\frac{10\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(8\gamma T/\delta)}.$
Proof.
For integer $K=\lceil\gamma T\rceil$. We first have that for $k\leq K$
$\tilde{y}_{k\lceil\frac{1}{\eta\gamma}\rceil}=(1-\eta\gamma)^{\lceil\frac{1}{\eta\gamma}\rceil}\tilde{y}_{(k-1)\lceil\frac{1}{\eta\gamma}\rceil}+\eta\sum_{\tau=0}^{\lceil\frac{1}{\eta\gamma}\rceil}(1-\eta\gamma)^{\lceil\frac{1}{\eta\gamma}\rceil-t}{g}_{(k-1)\lceil\frac{1}{\eta\gamma}\rceil+\tau}.$
Denote $\bar{g}_{k}=\eta\sum_{\tau=0}^{\lceil\frac{1}{\eta\gamma}\rceil}(1-\eta\gamma)^{\lceil\frac{1}{\eta\gamma}\rceil-t}{g}_{(k-1)\lceil\frac{1}{\eta\gamma}\rceil+\tau}$, then ${g}_{k}$ is a normal vector with variance
$\eta^{2}\sum_{\tau=0}^{\lceil\frac{1}{\eta\gamma}\rceil}(1-\eta\gamma)^{2(\lceil\frac{1}{\eta\gamma}\rceil-t)}\sigma^{2}\mathcal{I}\leq\frac{\eta\sigma^{2}}{2\gamma-\eta\gamma^{2}}\leq\frac{\eta\sigma^{2}}{\gamma}.$
Further, denote $Y_{k}=y_{k\lceil\frac{1}{\eta\gamma}\rceil}$ and $e_{\gamma}=(1-\eta\gamma)^{\lceil\frac{1}{\eta\gamma}\rceil}<\frac{1}{e}$, then
$Y_{k}=e_{\gamma}Y_{k-1}+\bar{g}_{k-1}=\sum_{i\leq k-1}e_{\gamma}^{i-1}{g}_{k-i}$
Then each variable $Y_{k}$ is also a Gaussian variable with variance smaller than
$\sum_{i\leq k-1}e_{\gamma}^{2(i-1)}\mathbb{E}[\bar{g}_{k-i}\bar{g}_{k-i}^{T}]\leq\frac{1}{1-1/e^{2}}\frac{\eta\sigma^{2}}{\gamma}\mathcal{I}\leq\frac{2\eta\sigma^{2}}{\gamma}\mathcal{I}.$
Hence, by Lemma A.35, for each $k$, it holds that
$\mathbb{P}(\Big{|}Y_{k}\Big{|}>\frac{2\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(4K/\delta)}))<\delta/2K.$
Using union bound,
$\mathbb{P}(\exists k\leq K,\Big{|}Y_{k}\Big{|}>\frac{2\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(4K/\delta)})<\delta/2.$
We now proceed to bound the distance of $y_{t}$ compared with close $Y_{k}$, without loss of generality, considering $k=0$, we will define a new process called $m_{t}$ satisfying that
$m_{t}=\sum_{k\leq t}(1-\eta\gamma)^{\lceil\frac{1}{\eta\gamma}\rceil-k}{g}_{k}.$
Then $m_{t}$ is a martingale and each $m_{t}$ is a Gaussian vector. In particular, $m_{\lceil\frac{1}{\eta\gamma}\rceil}=\bar{g}_{1}$. This further suggests that $\|m_{t}\|^{2}$ is a super martingale
$\mathbb{E}[\|m_{t}\|_{2}^{2}\mid m_{t-1}]\geq\|m_{t-1}\|_{2}^{2}.$
By Doob’s lemma (Lemma A.36)
$\mathbb{P}(\sup_{t\leq\lceil\frac{1}{\eta\gamma}\rceil}\|m_{t}\|_{2}^{2}>C^{2})$ $\leq\mathbb{P}(\sup_{t\leq\lceil\frac{1}{\eta\gamma}\rceil}\exp(\lambda\|m_{t}\|_{2}^{2})>\exp(\lambda C^{2}))$
$\leq\mathbb{E}[{\exp(\lambda\|m_{\lceil\frac{1}{\eta\gamma}\rceil}\|_{2}^{2}-\lambda C^{2})}]$
$=\mathbb{E}[{\exp(\lambda\|\bar{g}_{1}\|_{2}^{2}-\lambda C^{2})}].$
Following the same line of proof as Lemma A.35, we have that
$\mathbb{P}(\sup_{t\leq\lceil\frac{1}{\eta\gamma}\rceil}\|m_{t}\|_{2}>\frac{2\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(4K/\delta)})\leq\delta/2K$
We further note that $\Big{|}y_{t}-Y_{0}\Big{|}\leq(1-\eta\gamma)^{-\lceil\frac{1}{\eta\gamma}\rceil}m_{t}\leq 4m_{t}$. We have that for any $k<K$
$\mathbb{P}(\sup_{t\leq\lceil\frac{1}{\eta\gamma}\rceil}\|y_{t}-Y_{0}\|_{2}>\frac{8\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(4K/\delta)})\leq\delta/2K$
Combining with the bound on $Y_{k}$, we have that
$\mathbb{P}(\sup_{0\leq t\leq T}|y_{t}|>\frac{10\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(8\gamma T/\delta)}))\leq\delta.$
The proof is then complete. ∎
The following lemma states that $y(t)$ and $\tilde{y}(t)$ are close with high probability.
Lemma A.28.
Assume function $h(y)$ is $\gamma$-strong convex in $\mathcal{B}(0,r)$ and has a minimizer at 00,then for $\delta\in(0,1)$, under Assumption 6, it holds that with probability $1-\delta$,
$\forall t<T/\eta,\|\tilde{y}(t)-y(t)\|_{2}\leq 400{\eta\rho\sigma^{2}d\log(8\gamma T/\delta)}/{\gamma^{2}},$ $y(t)\in\mathcal{B}(0,r),\tilde{y}(t)\in\mathcal{B}(0,r)$
Proof.
By Lemma A.27, with probability $1-\delta$,
$\forall t<T/\eta,\|\tilde{y}_{t}\|_{2}^{2}\leq\frac{100{\eta}\sigma^{2}}{{\gamma}}{d\log(8\gamma T/\delta)}$
We will use $C$ as a shorthand for $100{d\log(8\gamma T/\delta)}$. Under such scenario, define $\nu_{t}=\tilde{y}(t)-y(t)$, we will prove by induction for $t\leq T/\eta$ that
$\|\nu_{t}\|_{2}\leq 4{\eta\rho\sigma^{2}C}/{\gamma^{2}},y_{t}\in\mathcal{B}(0,r).$ (28)
Clearly $\nu_{0}=0$, satisfies the induction hypothesis. Assuming Equation 28 hold for $t$, then
$y(t+1)$ $=y(t)-\eta\nabla L(y(t))-\eta{g}_{t}$
$=y(t)-\eta\nabla^{2}L(0)y(t)+e_{t}-\eta{g}_{t}$
$=\tilde{y}_{t}(1-\eta\nabla^{2}L(0))-\eta{g}_{t}+\nu_{t}(1-\eta\nabla^{2}L(0))+e_{t}.$
Here $\|e_{t}\|=\|-\eta(\nabla L(y(t))-\eta\nabla^{2}L(0)y(t))\|_{2}\leq\eta\rho\|y(t)\|_{2}^{2}\leq 2\eta\rho(\|\tilde{y}_{t}\|_{2}^{2}+\|\nu_{t}\|_{2}^{2})$. Hence we have that
$\|\nu_{t+1}\|_{2}\leq(1-\eta\gamma)\|\nu_{t}\|_{2}+2\eta\rho(\|\tilde{y}_{t}\|_{2}^{2}+\|\nu_{t}\|_{2}^{2}).$
As $\|\nu_{t}\|\leq 4\frac{\eta\rho\sigma^{2}C}{\gamma^{2}}\leq\frac{\gamma}{4\rho}$, we have that $22\rho\|\nu_{t}\|_{2}^{2}\leq\eta\gamma\|\nu_{t}\|_{2}^{2}/$ 2.
Hence
$\|\nu_{t+1}\|_{2}$ $\leq(1-\eta\gamma/2)\|\nu_{t}\|_{2}+2\eta\rho\frac{\eta\sigma^{2}C}{\gamma}$
$=(1-\eta\gamma/2)\|\nu_{t}\|_{2}+2\frac{\eta^{2}\rho\sigma^{2}C}{\gamma}$
By induction $\|\nu_{t}\|_{2}\leq 4{\eta\rho\sigma^{2}C}/{\gamma^{2}}$. It is then easy to check $\|\nu_{t+1}\|_{2}\leq 4{\eta\rho\sigma^{2}C}/{\gamma^{2}}$. ∎
The following lemma tracks the changes of $\mathbb{E}[h(y(t))]$.
Lemma A.29.
Assume function $h(y)$ is $\gamma$-strong convex in $\mathcal{B}(0,r)$ and has a minimizer at 00,then for $\delta\in(0,1)$, denote $100\log(8\gamma T/\delta)$ as $C$, under Assumption 6, it holds that $\forall t\in[1/\eta\gamma,T/\eta]$,
$\Big{|}\mathbb{E}[h(\tilde{y}_{t})]-\eta\sigma^{2}d/2\Big{|}\leq\eta^{2}\sigma^{2}\mathrm{Tr}(H)+\frac{\Delta\rho C}{\gamma^{2}}d\eta\sigma^{2}+\rho\left(\frac{d\eta\sigma^{2}C}{{\gamma}}\right)^{3/2}+2\delta M+\delta\eta\sigma^{2}d/2$
Proof.
By Lemma A.28, with probability $1-\delta$,
$\|\tilde{y}(t)-y(t)\|_{2}\leq 4{\eta\rho d\sigma^{2}C}/{\gamma^{2}},y(t)\in\mathcal{B}(0,r),\tilde{y}(t)\in\mathcal{B}(0,r).$
Define this event as $\mathcal{E}_{1}$.
Hence
$\Big{|}\mathbb{E}[h(y_{t})]-\mathbb{E}[h(\tilde{y}_{t})]\Big{|}$
$\leq$ $\Big{|}\mathbb{E}[h(\tilde{y}_{t})-h(y_{t})\mid\mathcal{E}_{1}]\mathbb{P}(\mathcal{E}_{1})+\mathbb{E}[h(\tilde{y}_{t})-h(y_{t})\mid\mathcal{E}_{1}^{c}]\mathbb{P}(\mathcal{E}_{1}^{c})\Big{|}$
$\leq$ $4\frac{\Delta\rho C}{\gamma^{2}}\eta d\sigma^{2}+\delta M.$
For $\|y\|_{2}<r$, it holds that
$\|h(y)-y^{T}\nabla^{2}h(0)y\|_{2}\leq\rho\|y\|_{2}^{3}.$
By Lemma A.27, with probability $1-\delta$, for $\eta<1/\gamma_{\mathrm{max}}$, it holds that for any $t\leq T/\eta$,
$\|\tilde{y}_{t}\|_{2}^{2}\leq\frac{d\eta\sigma^{2}C}{{\gamma}}<r^{2}.$
Define this event as $\mathcal{E}_{2}$.
Denote $H=\nabla^{2}h(0)$, we have that,
$\Big{|}\mathbb{E}[h(\tilde{y}_{t})]-\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}]\Big{|}$
$=$ $\Big{|}\mathbb{E}[h(\tilde{y}_{t})\mid\mathcal{E}_{2}]\mathbb{P}(\mathcal{E}_{2})-\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}]+\mathbb{E}[h(\tilde{y}_{t})\mid\mathcal{E}_{2}^{c}]\mathbb{P}(\mathcal{E}_{2}^{c})\Big{|}$
$\leq$ $\delta M+\Big{|}\mathbb{E}[h(\tilde{y}_{t})\mid\mathcal{E}_{2}]\mathbb{P}(\mathcal{E}_{2})-\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}]\Big{|}$
$\leq$ $\delta M+\rho\left(\frac{\eta d\sigma^{2}C}{{\gamma}}\right)^{3/2}+\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}\mid\mathcal{E}_{2}^{c}]$
Combining the both and we have that
$\Big{|}\mathbb{E}[h(y_{t})]-\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}]\Big{|}\leq 4\frac{\Delta\rho C}{\gamma^{2}}\eta d\sigma^{2}+\rho\left(\frac{\eta d\sigma^{2}C}{{\gamma}}\right)^{3/2}+2\delta M+\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}\mid\mathcal{E}_{2}^{c}].$
Here the covariance of $\tilde{y}_{t}$, denoted as $\Sigma_{t}$ satisfies that
$\Sigma_{t+1}=(\mathcal{I}-\eta H)^{2}\Sigma_{t}+\sigma^{2}\eta^{2}\mathcal{I}.$
Therefore
$\Sigma_{t}-\eta\sigma^{2}(2\eta H-\eta^{2}H^{2})^{-1}=(\mathcal{I}-\eta H)^{2}(\Sigma_{t-1}-\eta\sigma^{2}(2\eta H-\eta^{2}H^{2})^{-1}).$
$\Sigma_{t}=\sigma^{2}(2H-\eta H^{2})^{-1}(\mathcal{I}-(\mathcal{I}-\eta H)^{2t}).$
Hence assuming the eigenvalues of $H$ is $\gamma_{1},\dots,\gamma_{d}$
$\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}]=\mathrm{Tr}(\Sigma_{t}H)=\eta\sigma^{2}\sum_{i=1}^{d}\frac{1}{2-\eta\gamma_{i}}(1-(1-\eta\gamma_{i})^{2t}).$
When $t\geq\frac{1}{\eta\gamma_{i}},\eta\gamma_{i}<1/2$, it holds that
$\Big{|}\eta\sigma^{2}\frac{1}{2-\eta\gamma_{i}}(1-(1-\eta\gamma_{i})^{2t})-\eta\sigma^{2}/2\Big{|}$
$=$ $\eta\sigma^{2}\frac{(1-(1-\eta\gamma_{i})^{2t})}{2-\eta\gamma_{i}}-\eta\sigma^{2}/2$
$\leq$ $\eta\sigma^{2}\left(\frac{1}{2-\eta\gamma_{i}}-\frac{1}{2}\right)$
$\leq$ $\eta^{2}\sigma^{2}\gamma_{i}/2.$
Hence
$\Big{|}\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}]-d\eta\sigma^{2}/2\Big{|}\leq\mathrm{Tr}(H)\eta^{2}\sigma^{2}/2.$
Further, let $u_{t}=\Sigma_{t}^{-1/2}\tilde{y}_{t}$, under $\mathcal{E}_{2}^{c}$, we have that
$\|u_{t}\|_{2}^{2}\geq\lambda_{min}(\Sigma_{t}^{-1})\|\tilde{y}_{t}\|_{2}^{2}=\lambda_{min}\left((2H-\eta H^{2})(\mathcal{I}-(\mathcal{I}-\eta H)^{2t})^{-1}\right)\|\tilde{y}_{t}\|_{2}^{2}/\sigma^{2}\geq d\sigma^{2}C.$
As $u_{t}$ is isometric Gaussian,
$\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}\mid\mathcal{E}_{2}^{c}]$ $\leq\mathbb{E}[u_{t}^{T}(\Sigma_{t}^{1/2})^{T}H\Sigma_{t}^{1/2}u_{t}\mid\|u_{t}\|_{2}^{2}\geq d\sigma^{2}C]$
$=\mathbb{E}[u_{t}^{T}(\Sigma_{t}^{1/2})^{T}H\Sigma_{t}^{1/2}u_{t}]\frac{\mathbb{E}[\|u_{t}\|_{2}^{2}\mid\|u_{t}\|_{2}^{2}\geq d\sigma^{2}C]}{\mathbb{E}[\|u_{t}\|^{2}]}$
$\leq d\eta\sigma^{2}\frac{\mathbb{E}[\|u_{t}\|_{2}^{2}\mid\|u_{t}\|_{2}^{2}\geq d\sigma^{2}C]}{\mathbb{E}[\|u_{t}\|^{2}]}$
Plugging in the density function of $\|u_{t}\|_{2}$, we have that
[…]
Let $r^{\prime}=\sqrt{\frac{d}{d+1}}r$, then
[…]
Hence, we have that
$\mathbb{E}[(\tilde{y}_{t})^{T}H\tilde{y}_{t}\mid\mathcal{E}_{2}^{c}]\leq 4e^{-C/2}d\eta\sigma^{2}\leq\delta d\eta\sigma^{2}/2.$
Putting together, we have that,
$\Big{|}\mathbb{E}[h(\tilde{y}_{t})]-\eta\sigma^{2}d/2\Big{|}\leq\eta^{2}\sigma^{2}\mathrm{Tr}(H)+\frac{\Delta\rho C}{\gamma^{2}}d\eta\sigma^{2}+\rho\left(\frac{d\eta\sigma^{2}C}{{\gamma}}\right)^{3/2}+2\delta M+\delta\eta\sigma^{2}d/2.$
The proof is then complete. ∎
We will now state the complete version of Theorem 3.4.
Assumption 7 (Sufficient Small Learning Rate).
We will assume the following for constant $\delta\in(0,1]$ and learning rate $\eta$:
  1. 1.
    $\eta<1/(2\gamma_{\mathrm{max}})$.
  2. 2.
    $\eta\leq{\gamma^{3}}/{(1600\rho^{2}\sigma^{2}d\log(8\gamma T/\delta))}$.
  3. 3.
    $10\frac{\sqrt{\eta}\sigma}{\sqrt{\gamma}}\sqrt{d\log(8\gamma T/\delta)}+400\eta\rho\sigma^{2}d\log(8\gamma T/\delta)/\gamma^{2}\leq r$.
Theorem A.30 (Complete version of Theorem 3.4).
If a loss $L$ is a river valley (Definition 3.1) and satisfies Assumptions 3 and 4, for any constants $\delta\in(0,1)$ and $T>1/\gamma$, for sufficiently small learning rate $\eta$ satisfying Assumption 7, the iterate defined in Equation 4 with $\eta_{t}=\eta$, satisfies that for any integer $t\in[1/{\eta\gamma},T/\eta]$, there exists a $\tilde{T}$ satisfying that,
$\mathbb{E}[L\left(\tilde{w}(t)\right)]-L(x(\tilde{T}))$ $={(d-1)\eta\sigma^{2}}/{2}+\epsilon_{L}$
where $\epsilon_{t}=4\eta\gamma_{\mathrm{flat}}$ and $|\epsilon_{L}|\leq\tau\eta^{2}\sigma^{2}+\rho(C{d\eta\sigma^{2}}/{\gamma})^{3/2}+C\kappa^{\prime}d\eta\sigma^{2}+\delta(2M+\eta\sigma^{2}d)\ll(d-1)\eta\sigma^{2}$ with $C=200\log(64\gamma T/\delta)$.
Proof.
By Lemma A.26, we can write
$L(w)=h(w-\Phi(w))+L(\Phi(w)).$
Hence we can separate the dynamics of Equation 4 into two parts, namely $w=\Phi(w)+(w-\Phi(w))$. It is easy to check that when constrained on range of ${P}_{S}$, $h(y)$ satisfies Assumption 6. Hence, we can use Lemma A.29 to control $h(w_{t}-\Phi(w_{t}))$. For $\Phi(w_{t})$, the iterates is running a gradient descent with learning rate $\eta$ on $\mathcal{M}$ and we can use proof analogous to the proof of Theorem 3.3 to show that if $\Phi(w_{t})=x(\tilde{T}(t,\eta))$, then there exists $T_{0}$, such that
$\tilde{T}(t,\eta)\in[T_{0}+(1-4\eta\gamma_{\mathrm{flat}})\eta t,T_{0}+(1+4\eta\gamma_{\mathrm{flat}})\eta t].$
This completes the proof. ∎
A.7 Proof of Theorem 3.5
We will first state the complete version of Theorem 3.5.
Theorem A.31.
Under the setting of Theorem A.30, the SGD iterates (defined in Equation 4) with the learning rate schedule defined in Equation 5 satisfies that for any integer $t\in[t_{s},1.1t_{s}]$, there exists a $\tilde{T}\in[(1-\epsilon_{t})T(t),(1+\epsilon_{t})T(t)]$ satisfying that,
$\mathbb{E}[L\left(\tilde{w}(t)\right)]-L(x(\tilde{T}))$ $\leq{(d-1){\color[rgb]{1,0,0}{\eta_{t}}}\sigma^{2}}/{2}+\epsilon_{L}$
with $T(t)=T+\sum\limits_{k=t_{s}}^{t}\eta_{k}$.
Proof.
The proof is analogous to Theorem A.30 and Lemma A.29. We will omit the detail derivation and only focus on deriving the variance of corresponding $\tilde{y}_{t}$.
$\tilde{y}(t+1)=\tilde{y}(t)-\eta_{t}H\tilde{y}(t)-\eta_{t}{g}_{t},w(0)=0,\mathbf{{g}_{t}}\sim\mathcal{N}(0,\sigma^{2}\mathcal{I}),\tilde{y}(0)=0.$
Here the covariance of $\tilde{y}_{t}$, denoted as $\Sigma_{t}$ satisfies that
$\Sigma_{t+1}=(\mathcal{I}-\eta_{t}H)^{2}\Sigma_{t}+\sigma^{2}\eta_{t}^{2}\mathcal{I}.$
If we consider $i$-th eigenvector of $H$ as $v_{i}$, and denote $\sigma_{t,i}=v_{i}^{\top}\Sigma v_{i}$.
Analogous to the proof of Theorem A.30, $\Big{|}\sigma_{t_{s},i}-\frac{\eta\sigma^{2}}{\gamma_{i}}\Big{|}\leq\frac{4\eta^{2}\sigma^{2}}{\gamma_{i}}$.
We further have that
$\sigma_{t_{s}+r+1,i}$ $=(1-\frac{\eta}{2+r\eta\gamma}\gamma_{i})^{2}\sigma_{t_{s}+r,i}+\sigma^{2}\frac{\eta^{2}}{(2+r\eta\gamma)^{2}}.$
Then by induction, we can prove that for $r\geq 0$
$\sigma_{t_{s}+r+1,i}\leq\frac{\sigma^{2}}{\gamma_{i}}\frac{\eta}{2+r\eta\gamma}+\frac{4\eta^{2}\sigma^{2}}{\gamma_{i}}=\frac{\sigma^{2}}{\gamma_{i}}\eta_{t_{s}+r+1}+\frac{4\eta^{2}\sigma^{2}}{\gamma_{i}}.$
The rest follows the proof of Theorem A.30. ∎
A.8 Proof of Lemmas 4.1 and A.33
In this section, we will denote $\frac{\exp(\Theta_{i,j})}{\sum_{j=1}^{m}{\exp(\Theta_{i,j})}}$ as $\mathcal{Q}_{i,j}$
Lemma A.32.
The loss defined $L$ in Equation 6 satisfies that
$(\nabla L(\Theta))_{(i,j)}$ $=\mathcal{P}_{i,j}-\mathcal{Q}_{i,j}.$
$(\nabla^{2}L(\Theta))_{(i,j),(i^{\prime},j^{\prime})}$ $=\mathbf{1}(i=i^{\prime})(\mathcal{Q}_{i,j}\mathbf{1}(j=j^{\prime})-\mathcal{Q}_{i,j}\mathcal{Q}_{i,j^{\prime}}).$
Proof.
The loss satisfies that
$L(\Theta)=\sum_{i=1}^{n}(\sum_{j=1}^{m}\mathcal{P}_{i,j}\Theta_{i,j})-\log(\sum_{j^{\prime}=1}^{m}\mathcal{P}_{i,j^{\prime}}).$
Hence,
$(\nabla L(\Theta))_{(i,j)}=\mathcal{P}_{i,j}-\mathcal{Q}_{i,j}.$
Taking differentiation for another time yields the desired result. ∎
Proof of Lemma 4.1.
This can be done by directly summing diagonal entries in Lemma A.32. ∎
Assumption 8.
We will assume there exists constant $\gamma$ and positive integer $n^{\prime}<n$ such that $\mathcal{P}$ satisfies the following assumption,
  1. 1.
    For any $i\leq n^{\prime}$, $\forall j,\mathcal{P}_{i,j}>8\gamma$.
  2. 2.
    For any $i>n^{\prime}$, there exists $j_{i}$, $\mathcal{P}_{i,j^{\prime}}>1-\gamma$.
Theorem A.33.
Under Assumption 8, a generalized river with dimension $n^{\prime}m+(n-n^{\prime})$ exists in the loss landscape defined by $L$ in Equation 6.
Proof.
According to Lemma A.32, the Hessian for $L$ is block-diagonal. Now fixing a city $i$, we will analyze the eigenvalue distribution in this block. Let $q=[\mathcal{Q}_{i,j^{\prime}}]_{j^{\prime}\in[m]})$, then this block is $\mathrm{diag}(q)-qq^{T}$.
For all non-zero eigenvalue $\lambda$ for this block, there exists $v$ such that
$\mathrm{diag}(q)v-q^{T}vq=\lambda v.$
Hence, we have that
$v_{j}=\frac{q_{j}q^{T}v}{q_{j}-\lambda}$
This implies $\sum_{j=1}^{m}\frac{q_{j}^{2}}{q_{j}-c}=1$. We then have $\lambda\geq 0$ and there exists only one eigenvector corresponding to $\lambda=0$. For the rest nonzero eigenvalue, we have that $\lambda>\min q_{i}$.
Now if we consider the manifold $\mathcal{M}$ defined as
=$\mathcal{M}=\{\Theta\mid\forall i\leq n^{\prime},\mathcal{Q}_{i,j}=\mathcal{P}_{i,j};\forall i\geq n^{\prime},\mathcal{Q}_{i,j_{i}}>1-\gamma\}.$
Then for all $\Theta\in\mathcal{M}$, we have that the gradient is zero for all dimensions $(i,j)$ with $i\leq n^{\prime}$. Further, we know all the nonzero eigenvalues for these dimensions are at least $8\gamma$ by Assumption 8. For the rest of dimensions $(i,j)$ with $i>n^{\prime}$, by Lemma 4.1, the largest eigenvalue is bounded by $1-(1-\gamma)^{2}<2\gamma$. This shows that the gradient falls in the eigenspace spanned by the last $n^{\prime}m+(n-n^{\prime})$ eigenvectors, which concludes the proof. ∎
A.9 Technical Lemma
Lemma A.34.
If a function $F(t)$ satisfies that
$\frac{dF(t)}{dt}\leq-AF(t),$
then $F(t)\leq e^{-At}F(0)$.
Proof of Lemma A.34.
Consider $G(t)=F(t)e^{At}$, then
$dG(t)=e^{At}dF(t)+Ae^{At}F(t)\leq 0.$
Hence $G(t)\leq G(0)$. ∎
Lemma A.35.
If a random vector $g\sim\mathcal{N}(0,\Sigma)$ and $\delta\in(0,1)$, then it holds that
$\mathbb{P}(\|g\|_{2}\geq 2\sqrt{\mathrm{Tr}(\Sigma)}\sqrt{\log(2/\delta)})\leq\delta$
Proof of Lemma A.35.
Assume $\Sigma=Q\Lambda^{2}Q^{T}$ with $Q$ being an orthonormal matrix and $\Lambda$ being diagonal with diagonal $\lambda_{i}$ for $i\in[d]$, further let $g^{\prime}$ being a standard gaussian random vector, then $g$ follows the same distribution as that of $\Lambda Q^{T}g^{\prime}$, which is further identical to $\Lambda g^{\prime}$.
$\mathbb{P}(\|g\|_{2}\geq C)$
$=$ $\mathbb{P}(\|\Lambda g^{\prime}\|_{2}\geq C)$
$=$ $\mathbb{P}(\sum_{i=1}^{d}\Lambda_{i}^{2}(g^{\prime}_{i})^{2}\geq C^{2})$
$\leq$ $\mathbb{E}[\exp(t\sum_{i=1}^{d}\Lambda_{i}^{2}(g^{\prime}_{i})^{2}-tC^{2})].$
It is well known that the moment-generating function of $(g^{\prime}_{i})^{2}$ is
$E[\exp(t\Lambda_{i}^{2}(g^{\prime}_{i})^{2})]=\frac{1}{\sqrt{1-2t\Lambda_{i}^{2}}}.$
Hence
$\mathbb{P}(\|g\|_{2}\geq C)\leq$ $e^{-tC^{2}}\prod_{i=1}^{d}\frac{1}{\sqrt{1-2t\Lambda_{i}^{2}}}$
$\leq$ $\frac{e^{-tC^{2}}}{\sqrt{1-2t\mathrm{Tr}(\Sigma)}}.$
With $t=\frac{1}{4\mathrm{Tr}(\Sigma)}$, it holds that
$\mathbb{P}(\|g\|_{2}\geq C)\leq$ $2e^{-\frac{C^{2}}{4\mathrm{Tr}(\Sigma)}}.$
This concludes the proof. ∎
Lemma A.36 (Doob’s Inequality).
Let $X_{1},\dots,X_{n}$ as a positive submartingale adapted to filtration $\mathcal{F}_{1},\ldots,\mathcal{F}_{n}$, which means $X_{i}\leq\mathbb{E}[X_{i+1}\mid\mathcal{F}_{i}]$, then
$\mathbb{P}(\sup_{i\leq n}X_{i}>C)\leq\frac{\mathbb{E}[X_{n}]}{C}.$
Appendix B Omitted Experiments Details
We train LLaMA models with 4 parameter sizes using the Levanter framework for our study on WSD-S. For our theoretical study, we pretrain a 124M GPT-2 using the nanoGPT framework with a learning rate 6e-4 and train it with a batch size of 0.5M for 100k steps with warmup steps of 2k.
We hereby provide all the hyperparameters we used for the LLaMA and GPT-2 models training.
Model Hidden Dim Intermediate Dim Num Layers Num Heads Peak LR
0.1B LLaMa 768 3072 12 12 6e-4
0.3B LLaMa 1024 2048 24 16 6e-4
0.6B LLaMa 1536 6144 24 32 4e-4
1.2B LLaMa 2048 8096 16 32 4e-4
0.1B GPT-2 768 3072 12 12 6e-4
Table 1: Specifications for Different Sizes of LLaMa Models
We decay the model for the last $10\%$ of the training runs with one exception for 0.3B model using WSD method near 25k steps to avoid loss spikes. We outline the decayping and resuming point (the unit is 1k steps) we choose here:
Model 1st Decay Starts/Resume 2nd Decay Starts/Resume 3rd Decay Starts
0.1B LLaMa 11.25 / 12.5 22.5 / 25 48.75/ 53.75
0.3B LLaMa 11.25 / 12.5 22.5 / 25 48.75/ 53.75
0.6B LLaMa 11.25 / 12.5 22.5 / 25 48.75/ 53.75
1.2B LLaMa 11.25 /12.5 22.5 / 25 48.75/ 53.75
Table 2: Specifications for Decayping Steps for WSD-S Method
Model 1st Decay Starts/Ends 2nd Decay Starts/Ends 3rd Decay Starts Total Steps
0.1B LLaMa 11.25 / 12.5 22.5 / 25 45/ 50 53.75
0.3B LLaMa 11.25 / 12.5 22 / 25 45/ 50 54
0.6B LLaMa 11.25 / 12.5 22.5 / 25 45/ 50 53.75
1.2B LLaMa 11.25 /12.5 22.5 / 25 45/ 50 53.75
Table 3: Specifications for Decayping Steps for WSD Method